### v2

# Fourier integral operators on Orlicz modulation spaces

Abstract.
We establish continuity, compactness and Schatten-von Neumann properties for Fourier integral operators with amplitudes in Orlicz modulation spaces, when acting on other Orlicz modulation spaces themselves. The phase functions are non smooth and admit second order derivatives in suitable classes of modulation spaces.
0. Introduction
The aim of the paper is to investigate continuity and compactness properties for Fourier integral operators with non-smooth amplitudes (or symbols), when acting on Orlicz modulation spaces. In particular we extend continuity and compactness properties of the Fourier integral operators in [7, 8, 26, 60], where related results were obtained for such operators when acting on classical modulation spaces (of Lebesgue types). These earlier results are in turn extensions and generalizations of pioneering results and investigations performed by Boulkhemair in [5], where $L^{2}$-continuity of subclasses of our Fourier integral operators were established.
We recall that Orlicz versions of Lebesgue type spaces are obtained by replacing involved Lebesgue norms with Orlicz norms, which are parameterized with so-called Young functions. By choosing these Young functions in suitable ways one recovers the definitions of Lebesgue norms. Hence Orlicz type spaces extend the notions of analogous Lebesgue type spaces. In our situation, the family of Orlicz modulation spaces contains the corresponding family of classical modulation spaces.
A Fourier integral operator is an operator $\operatorname{Op}_{\varphi}(a)$, which is parameterized by the amplitude (or symbol) $a$ and phase function $\varphi$, defined on (suitable extensions of) the phase space. For appropriate $a\in\mathscr{S}^{\prime}(\mathbf{R}^{2d+m})$ and real-valued $\varphi\in C(\mathbf{R}^{2d+m})$, $\operatorname{Op}_{\varphi}(a)$ is an operator from $\mathscr{S}(\mathbf{R}^{d})$ to $\mathscr{S}^{\prime}(\mathbf{R}^{d})$, which is commonly defined as
$\operatorname{Op}_{\varphi}(a)f(x)$ $=(2\pi)^{-\frac{1}{2}(d+m)}\iint_{\mathbf{R}^{d+m}}a(x,y,\zeta)f(y)e^{i\varphi(x,y,\zeta)}\,dyd\zeta,$ (0.1)
$f$ $\in\mathscr{S}(\mathbf{R}^{d}),\quad x\in\mathbf{R}^{d}.$
(See [35] and Section 1 for notations.) To some extent, the amplitude $a$ quantifies the amplification, while the phase function $\varphi$ is linked to (non-homogeneous) modulations within the systems.
In various situations were Fourier integral operators are used, the right choice of the phase function $\varphi$ is crucial. In hyperbolic problems, which is a common field for applying Fourier integral operators, one usually assumes that $\varphi(x,y,\zeta)$ is smooth when $\zeta\neq 0$ and positively homogeneous of order one with respect to the $\zeta$ variable. In these hyperbolic situations, the adapted $\varphi(x,y,\zeta)$ usually fails to be differentiable in the $\zeta$ variable when $\zeta=0$. Furthermore, one usually impose strong regularity assumptions on the amplitudes, e. g. that they should belong to subclasses of the Hörmander class $S^{r}_{0,0}$, or that they should belong to so-called SG-classes. (See e. g. [4, 6, 17, 18, 19, 20, 45, 48, 49] and the references therein.)
Roughly speaking, in contrast to the previous assumptions adapted to hyperbolic problems, in our situation we use the framework in [5] by Boulkhemair and in [60] concerning (lack of) regularity assumptions on the amplitude $a$ and phase function $\varphi$. Especially we impose less restrictions on $\varphi(x,y,\zeta)$ outside $\zeta=0$, but stronger regularity with respect to $\zeta$ when $\zeta=0$, compared to the situation of hyperbolic problems. Furthermore, we stress that in similar ways as in [5, 60], the regularity assumptions on the amplitude $a$ are also relaxed compared to what is common for Fourier integral operators.
In order to be more specific, suppose that $\omega$, $\omega_{1}$, $\omega_{2}$ and $v$ are suitable weight functions, and that $\Phi$ is a suitable Young function. Also suppose that $a$ belongs to the modulation space $M^{\infty,1}_{(\omega)}(\mathbf{R}^{2d+m})$ (of so-called Sjöstrand class), and that
$\varphi^{\prime\prime}\in M^{\infty,1}_{(v)}(\mathbf{R}^{2d+m})$
satisfies the non-degeneracy condition
$\left|\det\left(\begin{matrix}\varphi^{\prime\prime}_{y,x}&&\varphi^{\prime\prime}_{\zeta,x}\\[4.30554pt] \varphi^{\prime\prime}_{y,\zeta}&&\varphi^{\prime\prime}_{\zeta,\zeta}\end{matrix}\right)\right|\geq d\hskip-0.80002pt\bar{}\hskip 1.99997pt.$ (0.2)
for some $d\hskip-0.80002pt\bar{}\hskip 1.99997pt>0$. Then Theorem 2.3 in Section 2 asserts that $\operatorname{Op}_{\varphi}(a)$ is continuous between the Orlicz modulation spaces $M^{\Phi}_{(\omega_{1})}(\mathbf{R}^{d})$ and $M^{\Phi}_{(\omega_{2})}(\mathbf{R}^{d})$. That is, the map
$\operatorname{Op}_{\varphi}(a):M^{\Phi}_{(\omega_{1})}(\mathbf{R}^{d})\to M^{\Phi}_{(\omega_{2})}(\mathbf{R}^{d})$ (0.3)
is continuous. If, more restrictive, $a$ is chosen in the completion $M^{\sharp,1}_{(\omega)}(\mathbf{R}^{2d+m})$ of $M^{1,1}_{(\omega)}(\mathbf{R}^{2d+m})$ in $M^{\infty,1}_{(\omega)}(\mathbf{R}^{2d+m})$, then the map (0.3) is compact. (See Theorem 2.4.)
We notice that we may choose $\omega$ such that $M^{\infty,1}_{(\omega)}$ contain $S^{r}_{0,0}$. In particular, as announced above, we put less restrictions on the amplitudes compared to the assumptions above adapted to hyperbolic problems.
We remark that our results extend and generalize related results in [60] in especially two different ways.
Firstly, we may choose the involved Young functions such that $M^{\Phi}_{(\omega_{j})}$ in (0.3) is equal to the classical modulation space $M^{p}_{(\omega_{j})}$, for any $p\in[1,\infty]$, $j=1,2$.
Secondly, in (0.3) we allow $\omega$, $\omega_{j}$ and $v$ to belong to the large class $\mathscr{P}_{s}$ of weights which are moderate by subexponential functions of degree $s^{-1}<1$, while in [60] it is required that the involed weights should belong to the smaller class $\mathscr{P}$ of polynomially moderate weights. By imposing these two significant restrictions, then our Theorem 2.3 in Section 2, as well as (0.3), essentially takes the form [60, Theorem 2.2].
In similar ways, we deduce several other continuity properties, including detailed compactness properties, for Fourier integral operators on Orlicz modulation spaces, which cover the continuity results in [5, 7, 8, 60]. For example, we investigate $\operatorname{Op}_{\varphi}(a)$ when the amplitude $a$ satisfies conditions which are rather similar to norm estimates with respect to the Orlicz modulation space $M^{\Phi}_{(\omega)}(\mathbf{R}^{2d+m})$. For suitable non-degeneracy conditions on $\varphi$, different compared to (0.2), we show that
$\operatorname{Op}_{\varphi}(a):M^{\Phi^{*}}_{(\omega_{1})}(\mathbf{R}^{d})\to M^{\Phi}_{(\omega_{2})}(\mathbf{R}^{d})$
is continuous. (See Theorem 2.6 in Section 2.)
In Section 3 we perform detailed studies on compactness for Fourier integral operators. Here we find necessary conditions on the amplitudes in order for corresponding Fourier integral operators should belong to certain Orlicz Schatten-von Neumann classes. We impose certain restrictions on the amplitudes. In the first step we assume that the amplitudes in (0.1) are independent of the $y$ variable. That is, they are of the form
$\operatorname{Op}_{\varphi}(a)f(x)$ $=(2\pi)^{-d}\iint_{\mathbf{R}^{2d}}a(x,\zeta)f(y)e^{i\varphi(x,y,\zeta)}\,dyd\zeta,$ (0.4)
$f$ $\in\mathscr{S}(\mathbf{R}^{d}),\quad x\in\mathbf{R}^{d}.$
Thereafter we consider a more general family of Fourier integral operators, given by
$\operatorname{Op}_{A,\varphi}(a)f(x)$ $=(2\pi)^{-d}\iint_{\mathbf{R}^{2d}}a(x-A(x-y),\zeta)f(y)e^{i\varphi(x,y,\zeta)}\,dyd\zeta,$ (0.5)
$f$ $\in\mathscr{S}(\mathbf{R}^{d}),\quad x\in\mathbf{R}^{d}.$
Again we deduce Orlicz Schatten-von Neumann properties for such Fourier integral operators. In fact, for amplitudes $a$ in (0.4) (or more general (0.5)), in suitable weighted $M^{\Phi}$ classes, and suitable phase functions $\varphi$, we show that corresponding Fourier integral operators are Schatten-von Neumann operators of order $\Phi$ from $M^{2}_{(\omega_{1})}$ to $M^{2}_{(\omega_{2})}$.
Finally we remark that by choosing
$m=d\quad\text{and}\quad\varphi(x,y,\zeta)\equiv\langle x-y,\zeta\rangle,$
then our Fourier integral operator (0.1) becomes the pseudo-differential operator
$\operatorname{Op}(a)f(x)=(2\pi)^{-d}\iint_{\mathbf{R}^{d}}a(x,y,\zeta)f(y)e^{i\langle x-y,\zeta\rangle}\,dyd\zeta.$ (0.6)
In particular, pseudo-differential operators are special cases of Fourier integral operators. Furthermore, if $A$ is a fixed real $d\times d$ matrix, and instead $a\in\mathscr{S}^{\prime}(\mathbf{R}^{2d})$ is an appropriate function or distribution on $\mathbf{R}^{2d}$ instead of $\mathbf{R}^{3d}$, then pseudo-differential operators of the form
$\operatorname{Op}_{A}(a)f(x)=(2\pi)^{-d}\iint_{\mathbf{R}^{2d}}a(x-A(x-y),\xi)f(y)e^{i\langle x-y,\xi\rangle}\,dyd\xi,$ (0.7)
can be considered as special case of operators in (0.5), as well as in (0.6).
On the other hand, by Fourier inversion formula and kernel theorems it follows that any continuous operator from $\mathscr{S}(\mathbf{R}^{d})$ to $\mathscr{S}^{\prime}(\mathbf{R}^{d})$ is given by (0.7) for a suitable choice of $a$. Consequently, the set of operators are in general not increased by passing from the more restricted formulation (0.7) to the more general formulations (0.1) via (0.5) or (0.6).
The additional assumptions on the phase function for the pseudo-differential operators in (0.6) and (0.7) lead to more general continuity properties, compared to what is possible for general Fourier integral operators in (0.1). For related continuity properties for pseudo-differential operators when acting on Orclicz modulation spaces, see [33, 62]. For some further extensions to more general modulation spaces, see [31, 61] and the references therein. For earlier approaches restricted to classical modulation spaces, see e. g. [10, 16, 56, 57] and the references therein.
1. Preliminaries
In the section we recall some basic facts on Gelfand-Shilov spaces, Orlicz spaces, Orlicz modulation spaces, pseudo-differential operators and Wigner distributions. We also give some examples on Young functions, Orlicz spaces and Orlicz modulation spaces. (See Examples 1.9 and 1.16.) Notice that Young functions are fundamental in the definition of Orlicz spaces and Orlicz modulation spaces).
1.1. Gelfand-Shilov spaces
For a real number $s>0$, the (standard Fourier invariant) Gelfand-Shilov space $\mathcal{S}_{s}(\mathbf{R}^{d})$ ($\Sigma_{s}(\mathbf{R}^{d})$) of Roumieu type (Beurling type) consists of all $f\in C^{\infty}(\mathbf{R}^{d})$ such that
$\|f\|_{\mathcal{S}_{s,h}}\equiv\sup_{\begin{subarray}{c}\alpha,\beta\in\mathbf{N}^{d}\\ x\in\mathbf{R}^{d}\end{subarray}}\frac{|x^{\beta}\partial^{\alpha}f(x)|}{h^{|\alpha+\beta|}(\alpha!\beta!)^{s}}$ (1.1)
is finite for some $h>0$ (for every $h>0$). We equip $\mathcal{S}_{s}(\mathbf{R}^{d})$ ($\Sigma_{s}(\mathbf{R}^{d})$) by the canonical inductive limit topology (projective limit topology) with respect to $h>0$, induced by the semi-norms defined in (1.1).
We have
$\mathcal{S}_{s}(\mathbf{R}^{d})$ $\hookrightarrow\Sigma_{t}(\mathbf{R}^{d})\hookrightarrow\mathcal{S}_{t}(\mathbf{R}^{d})\hookrightarrow\mathscr{S}(\mathbf{R}^{d})$ (1.2)
$\hookrightarrow\mathscr{S}^{\prime}(\mathbf{R}^{d})\hookrightarrow\mathcal{S}_{t}^{\prime}(\mathbf{R}^{d})\hookrightarrow\Sigma_{t}^{\prime}(\mathbf{R}^{d})\hookrightarrow\mathcal{S}_{s}^{\prime}(\mathbf{R}^{d}),\quad\frac{1}{2}\leq s<t,$
with dense embeddings. Here $A\hookrightarrow B$ means that the topological space $A$ is continuously embedded in the topological space $B$. We also have
$\mathcal{S}_{s}(\mathbf{R}^{d})=\Sigma_{t}(\mathbf{R}^{d})=\{0\},\qquad s<\frac{1}{2},\ t\leq\frac{1}{2}.$
The Gelfand-Shilov distribution spaces $\mathcal{S}_{s}^{\prime}(\mathbf{R}^{d})$ and $\Sigma_{s}^{\prime}(\mathbf{R}^{d})$, of Roumieu and Beurling types respectively, are the (strong) duals of $\mathcal{S}_{s}(\mathbf{R}^{d})$ and $\Sigma_{s}(\mathbf{R}^{d})$, respectively. It follows that if $\mathcal{S}_{s,h}^{\prime}(\mathbf{R}^{d})$ is the $L^{2}$-dual of $\mathcal{S}_{s,h}(\mathbf{R}^{d})$ and $s\geq\frac{1}{2}$ ($s>\frac{1}{2}$), then $\mathcal{S}_{s}^{\prime}(\mathbf{R}^{d})$ ($\Sigma_{s}^{\prime}(\mathbf{R}^{d})$) can be identified with the projective limit (inductive limit) of $\mathcal{S}_{s,h}^{\prime}(\mathbf{R}^{d})$ with respect to $h>0$. It follows that
$\mathcal{S}_{s}^{\prime}(\mathbf{R}^{d})=\bigcap_{h>0}\mathcal{S}_{s,h}^{\prime}(\mathbf{R}^{d})\quad\text{and}\quad\Sigma_{s}^{\prime}(\mathbf{R}^{d})=\bigcup_{h>0}\mathcal{S}_{s,h}^{\prime}(\mathbf{R}^{d})$ (1.3)
for such choices of $s$ and $\sigma$, see [27, 40, 41] for details.
We let the Fourier transform $\mathscr{F}$ be given by
$(\mathscr{F}f)(\xi)=\widehat{f}(\xi)\equiv(2\pi)^{-\frac{d}{2}}\int_{\mathbf{R}^{d}}f(x)e^{-i\langle x,\xi\rangle}\,dx,\quad\xi\in\mathbf{R}^{d},$
when $f\in L^{1}(\mathbf{R}^{d})$. Here $\langle\,\cdot\,$$\cdot\,\rangle$ denotes the usual scalar product on $\mathbf{R}^{d}$. The Fourier transform $\mathscr{F}$ extends uniquely to homeomorphisms on $\mathscr{S}^{\prime}(\mathbf{R}^{d})$, $\mathcal{S}_{s}^{\prime}(\mathbf{R}^{d})$ and on $\Sigma_{s}^{\prime}(\mathbf{R}^{d})$. Furthermore, $\mathscr{F}$ restricts to homeomorphisms on $\mathscr{S}(\mathbf{R}^{d})$, $\mathcal{S}_{s}(\mathbf{R}^{d})$ and on $\Sigma_{s}(\mathbf{R}^{d})$, and to a unitary operator on $L^{2}(\mathbf{R}^{d})$. Similar facts hold true with partial Fourier transforms in place of Fourier transform.
Let $\phi\in\mathscr{S}(\mathbf{R}^{d})$ be fixed. Then the short-time Fourier transform $V_{\phi}f$ of $f\in\mathscr{S}^{\prime}(\mathbf{R}^{d})$ with respect to the window function $\phi$ is the tempered distribution on $\mathbf{R}^{2d}$, defined by
$V_{\phi}f(x,\xi)$ $=\mathscr{F}(f\,\overline{\phi(\,\cdot\,-x)})(\xi),\quad x,\xi\in\mathbf{R}^{d}.$ (1.4)
In some situations it is convenient to use the small modification
$T_{\phi}f(x,\xi)$ $=\mathscr{F}(f((\,\cdot\,+x))\,\overline{\phi})(\xi),\quad x,\xi\in\mathbf{R}^{d}$ (1.5)
of $V_{\phi}f$. By a straight-forward change of variables it follows that
$T_{\phi}f(x,\xi)=e^{i\langle x,\xi\rangle}V_{\phi}f(x,\xi)$
If $f,\phi\in\mathscr{S}(\mathbf{R}^{d})$, then it follows that
$V_{\phi}f(x,\xi)=(2\pi)^{-\frac{d}{2}}\int_{\mathbf{R}^{d}}f(y)\overline{\phi(y-x)}e^{-i\langle y,\xi\rangle}\,dy,\quad x,\xi\in\mathbf{R}^{d}.$
By [58, Theorem 2.3] it follows that the definition of the map $(f,\phi)\mapsto V_{\phi}f$ from $\mathscr{S}(\mathbf{R}^{d})\times\mathscr{S}(\mathbf{R}^{d})$ to $\mathscr{S}(\mathbf{R}^{2d})$ is uniquely extendable to a continuous map from $\mathcal{S}_{s}^{\prime}(\mathbf{R}^{d})\times\mathcal{S}_{s}^{\prime}(\mathbf{R}^{d})$ to $\mathcal{S}_{s}^{\prime}(\mathbf{R}^{2d})$, and restricts to a continuous map from $\mathcal{S}_{s}(\mathbf{R}^{d})\times\mathcal{S}_{s}(\mathbf{R}^{d})$ to $\mathcal{S}_{s}(\mathbf{R}^{2d})$. The same conclusion holds with $\Sigma_{s}$ in place of $\mathcal{S}_{s}$, at each occurrence.
In the following proposition we give characterizations of Gelfand-Shilov spaces and their distribution spaces in terms of estimates of the short-time Fourier transform. We omit the proof since the first part follows from [32, Theorem 2.7] and the second part from [58, Theorem 2.5]. See also [15] for related results. Here and in what follows, the notation $A(\theta)\lesssim B(\theta)$, $\theta\in\Omega$, means that there is a constant $c>0$ such that $A(\theta)\leq cB(\theta)$ holds for all $\theta\in\Omega$. We also set $A(\theta)\asymp B(\theta)$ when $A(\theta)\lesssim B(\theta)\lesssim A(\theta)$.
Proposition 1.1.
Let $s\geq\frac{1}{2}$ ($s>\frac{1}{2}$), $\phi\in\mathcal{S}_{s}(\mathbf{R}^{d})\setminus 0$ ($\phi\in\Sigma_{s}(\mathbf{R}^{d})\setminus 0$) and let $f$ be a Gelfand-Shilov distribution on $\mathbf{R}^{d}$. Then the following is true:
  1. (1)
    $f\in\mathcal{S}_{s}(\mathbf{R}^{d})$ ($f\in\Sigma_{s}(\mathbf{R}^{d})$), if and only if
    $|V_{\phi}f(x,\xi)|\lesssim e^{-r(|x|^{\frac{1}{s}}+|\xi|^{\frac{1}{s}})},\quad x,\xi\in\mathbf{R}^{d},$ (1.6)
    for some $r>0$ (for every $r>0$).
  2. (2)
    $f\in\mathcal{S}_{s}^{\prime}(\mathbf{R}^{d})$ ($f\in\Sigma_{s}^{\prime}(\mathbf{R}^{d})$), if and only if
    $|V_{\phi}f(x,\xi)|\lesssim e^{r(|x|^{\frac{1}{s}}+|\xi|^{\frac{1}{s}})},\quad x,\xi\in\mathbf{R}^{d},$ (1.7)
    for every $r>0$ (for some $r>0$).
In our investigations, also compactly supported elements in Gelfand-Shilov spaces appears. For this reason we set
$\mathcal{D}_{s}(K)\equiv\mathcal{S}_{s}(\mathbf{R}^{d})\bigcap\mathscr{E}^{\prime}(K)\quad\text{and}\quad\mathcal{D}_{0,s}(K)\equiv\Sigma_{s}(\mathbf{R}^{d})\bigcap\mathscr{E}^{\prime}(K),$
when $K\subseteq\mathbf{R}^{d}$ is compact, with topologies induced by the topologies from $\mathcal{S}_{s}(\mathbf{R}^{d})$ and $\Sigma_{s}(\mathbf{R}^{d})$, respectively. We also let
$\mathcal{D}_{s}(\mathbf{R}^{d})=\bigcup_{j=1}^{\infty}\mathcal{D}_{s}(K_{j})\quad\text{and}\quad\mathcal{D}_{0,s}(\mathbf{R}^{d})=\bigcup_{j=1}^{\infty}\mathcal{D}_{0,s}(K_{j}),$ (1.8)
where
$K_{j}=\{\,x\in\mathbf{R}^{d}\,;\,|x|\leq j\,\}$
is the closed ball of radius $j$ with center at origin. We let the topologies of $\mathcal{D}_{s}(\mathbf{R}^{d})$ and $\mathcal{D}_{0,s}(\mathbf{R}^{d})$ be the inductive limit topologies of $\mathcal{D}_{s}(K_{j})$ and $\mathcal{D}_{0,s}(K_{j})$ with respect to $j$.
It is well-known that if $s>1$, then
$\mathcal{D}_{0,s}(\mathbf{R}^{d})\subseteq\mathcal{D}_{s}(\mathbf{R}^{d})\subseteq C_{0}^{\infty}(\mathbf{R}^{d})\subseteq\mathscr{S}(\mathbf{R}^{d}),$
with dense embeddings. (See e. g. Sections 1.3 and 8.4 in [35].) On the other hand, if $s\leq 1$, then
$\mathcal{D}_{0,s}(\mathbf{R}^{d})=\mathcal{D}_{s}(\mathbf{R}^{d})=\{0\},$
that is, the spaces in (1.8) becomes trivial. This follows from the fact that for $s\leq 1$, then all elements in spaces in (1.8) are real analytic functions.
1.2. Weight functions
A weight or weight function on $\mathbf{R}^{d}$ is a positive function $\omega\in L^{\infty}_{loc}(\mathbf{R}^{d})$ such that $1/\omega\in L^{\infty}_{loc}(\mathbf{R}^{d})$. The weight $\omega$ is called moderate, if there is a positive weight $v$ on $\mathbf{R}^{d}$ such that
$\omega(x+y)\lesssim\omega(x)v(y),\qquad x,y\in\mathbf{R}^{d}.$ (1.9)
If $\omega$ and $v$ are weights on $\mathbf{R}^{d}$ such that (1.9) holds, then $\omega$ is also called $v$-moderate. We note that (1.9) implies that $\omega$ fulfills the estimates
$v(-x)^{-1}\lesssim\omega(x)\lesssim v(x),\quad x\in\mathbf{R}^{d}.$ (1.10)
We let $\mathscr{P}_{E}(\mathbf{R}^{d})$ be the set of all moderate weights on $\mathbf{R}^{d}$.
It can be proved that if $\omega\in\mathscr{P}_{E}(\mathbf{R}^{d})$, then $\omega$ is $v$-moderate for some $v(x)=e^{r|x|}$, provided the positive constant $r$ is large enough (cf. [29]). That is, (1.9) implies
$\omega(x+y)\lesssim\omega(x)e^{r|y|}$ (1.11)
for some $r>0$. In particular, (1.10) shows that for any $\omega\in\mathscr{P}_{E}(\mathbf{R}^{d})$, there is a constant $r>0$ such that
$e^{-r|x|}\lesssim\omega(x)\lesssim e^{r|x|},\quad x\in\mathbf{R}^{d}.$ (1.12)
We say that $v$ is submultiplicative if $v$ is even and (1.9) holds with $\omega=v$. In the sequel, $v$ and $v_{j}$ for $j\geq 0$, always stand for submultiplicative weights if nothing else is stated.
For any $s>0$, we let $\mathscr{P}_{s}(\mathbf{R}^{d})$ be the set of all weights $\omega$ on $\mathbf{R}^{d}$ such that
$\omega(x+y)\lesssim\omega(x)e^{r|y|^{\frac{1}{s}}}$ (1.13)
holds for some $r>0$. In the same manner the set $\mathscr{P}_{0,s}(\mathbf{R}^{d})$ consists of all weights $\omega$ on $\mathbf{R}^{d}$ such that (1.13) is true for every $r>0$. We also let $\mathscr{P}(\mathbf{R}^{d})$ be the set of all $\omega\in\mathscr{P}_{E}(\mathbf{R}^{d})$ such that
$\omega(x+y)\lesssim\omega(x)(1+|y|)^{r}$
for some $r>0$.
Evidently,
$\mathscr{P}(\mathbf{R}^{d})\subseteq\mathscr{P}_{s_{1}}(\mathbf{R}^{d})\subseteq\mathscr{P}_{0,s_{2}}(\mathbf{R}^{d})\subseteq\mathscr{P}_{s_{2}}(\mathbf{R}^{d}),\qquad 0<s_{1}<s_{2}.$
On the other hand, in view of (1.13) it follows that all weights in $\mathscr{P}_{0,s}(\mathbf{R}^{d})$ and $\mathscr{P}_{s}(\mathbf{R}^{d})$ are moderate. Hence (1.11) gives
$\mathscr{P}_{s_{1}}(\mathbf{R}^{d})=\mathscr{P}_{0,s_{2}}(\mathbf{R}^{d})=\mathscr{P}_{E}(\mathbf{R}^{d})\quad\text{when}\quad s_{1}\geq 1,\ s_{2}>1.$
1.3. Orlicz Spaces
We recall that a function $\Phi:[0,\infty]\to[0,\infty]$ is called convex if
$\Phi(s_{1}t_{1}+s_{2}t_{2})\leq s_{1}\Phi(t_{1})+s_{2}\Phi(t_{2}),$
when $s_{j},t_{j}\in\mathbf{R}$ satisfy $s_{j},t_{j}\geq 0$ and $s_{1}+s_{2}=1,\ j=1,2$.
Definition 1.2.
A function $\Phi_{0}$ from $0,\infty]$ to $0,\infty]$ is called a Young function if the following is true:
  1. (1)
    $\Phi_{0}$ is convex;
  2. (2)
    $\Phi_{0}(0)=0$;
  3. (3)
    $\lim\limits_{t\to\infty}\Phi_{0}(t)=\Phi_{0}(\infty)=\infty$.
We observe that $\Phi_{0}$ and $\Phi$ in Definition 1.2 might not be continuous, because we permit $\infty$ as function value. For example,
$\Phi(t)=\begin{cases}0,&\text{when}\ t\leq a\\[4.30554pt] \infty,&\text{when}\ t>a\end{cases}$
is convex but discontinuous at $t=a$.
It is clear that $\Phi_{0}$ and $\Phi$ in Definition 1.2 are non-decreasing, because if $0\leq t_{1}\leq t_{2}$ and $s\in[0,1]$ is chosen such that $t_{1}=st_{2}$ and $\Phi_{0}$ is the same as in Definition 1.2, then
$\Phi_{0}(t_{1})=\Phi_{0}(st_{2}+(1-s)0)\leq s\Phi_{0}(t_{2})+(1-s)\Phi_{0}(0)\leq\Phi_{0}(t_{2}),$
since $\Phi_{0}(0)=0$ and $s\in[0,1]$. Hence every Young function is increasing.
Definition 1.3.
Let $\Phi$ be a Young function and let $\omega_{0}\in\mathscr{P}_{E}(\mathbf{R}^{d})$. Then the Orlicz space $L^{\Phi}_{(\omega_{0})}(\mathbf{R}^{d})$ consists of all measurable functions $f:\mathbf{R}^{d}\to\mathbf{C}$ such that
$\|f\|_{L^{\Phi}_{(\omega_{0})}}\equiv\inf\left\{\,\lambda>0\,;\,\int_{\Omega}\Phi\left(\frac{|f(x)\cdot\omega_{0}(x)|}{\lambda}\right)\,dx\leq 1\,\right\}$
is finite. Here $f$ and $g$ in $L^{\Phi}_{(\omega_{0})}(\mathbf{R}^{d})$ are equivalent if $f=g$ a.e.
In most of our situations we assume that $\Phi$ and $\Phi_{j}$ above are Young functions. A few properties for Wigner distributions in Section 2 are deduced when $\Phi$ and $\Phi_{j}$ are allowed to be Young functions. The reader who is not interested of such general results may always assume that all Young functions should be Young functions.
It is well-known that if $\Phi$ in Definition 1.3 is a Young function, then the space $L^{\Phi}_{(\omega_{0})}(\mathbf{R}^{d})$ and $L^{\Phi_{1},\Phi_{2}}_{(\omega)}(\mathbf{R}^{2d})$ is a Banach spaces (see e. g. Theorem 3 of III.3.2 and Theorem 10 of III.3.3 in [43]).
We refer to [62, Lemma 1.18] for the proof of the following lemma.
Lemma 1.4.
Let $\Phi,\Phi_{j}$ be Young functions, $j=1,2$, $\omega_{0},v_{0}\in\mathscr{P}_{E}(\mathbf{R}^{d})$ and $\omega,v\in\mathscr{P}_{E}(\mathbf{R}^{d}d)$ be such that $\omega_{0}$ is $v_{0}$-moderate and $\omega$ is $v$-moderate. Then $L^{\Phi}_{(\omega_{0})}(\mathbf{R}^{d})$ are invariant under translations, and
$\|f(\,\cdot\,-x)\|_{L^{\Phi}_{(\omega_{0})}}\lesssim\|f\|_{L^{\Phi}_{(\omega_{0})}}v_{0}(x),\quad f\in L^{\Phi}_{(\omega_{0})}(\mathbf{R}^{d}),\ x\in\mathbf{R}^{d}.$
In most situations we assume that the Young functions should satisfy the $\Delta_{2}$-condition (near origin), whose definition is recalled as follows.
Definition 1.5.
Let $\Phi:[0,\infty]\to[0,\infty]$ be a Young function. Then $\Phi$ is said to satisfy the $\Delta_{2}$-condition if there exists a constant $C>0$ such that
$\Phi(2t)\leq C\Phi(t)$ (1.14)
for every $t\in[0,\infty]$. The Young function $\Phi$ is said to satisfy local $\Delta_{2}$-condition or $\Delta_{2}$-condition near origin, if there are constants $r>0$ and $C>0$ such that (1.14) holds when $t\in[0,r]$.
Remark 1.6.
Suppose that $\Phi:[0,\infty]\to[0,\infty]$ is a Young function which satisfies (1.14) when $t\in[0,r]$ for some constants $r>0$ and $C>0$. Then it follows by straight-forward arguments that there is a Young function $\Phi_{0}$ (of the same order) which satisfies the $\Delta_{2}$-condition (on the whole $0,\infty$), and such that $\Phi_{0}(t)=\Phi(t)$ when $t\in[0,r]$).
Several duality properties for Orlicz spaces can be described in terms of Orlicz spaces with respect to Young conjugates, given in the following definition.
Definition 1.7.
Let $\Phi$ be a Young function. Then the conjugate Young function $\Phi^{*}$ is given by
$\Phi^{*}(t)\equiv\begin{cases}{\displaystyle{\sup_{s\geq 0}(st-\Phi(s)),}}&\text{when}\ t\in[0,\infty),\\[8.61108pt] \infty,&\text{when}\ t=\infty.\end{cases}$ (1.15)
Remark 1.8.
Let $p\in[1,\infty]$, and set $\Phi_{[p]}(t)=t^{p}$ when $p\in(0,\infty)$, and
$\Phi_{[\infty]}(t)=\begin{cases}0,&t\leq 1,\\[4.30554pt] \infty,&t>1.\end{cases}$
Then $L^{\Phi_{[p]}}(\mathbf{R}^{d})$ and its norm is equal to the classical Lebesgue space $L^{p}(\mathbf{R}^{d})$ and its norm.
Moreover, suppose $p_{1},p_{2}\in[1,\infty]$ and let
$\Phi(t)=\begin{cases}{\frac{t^{p_{2}}}{p_{2}}},&0\leqslant t\leqslant 1,\\[4.30554pt] {\frac{t^{p_{1}}}{p_{1}}+\frac{1}{p_{2}}-\frac{1}{p_{1}}},&t>1.\end{cases}$
Here, we interpret $\frac{t^{\infty}}{\infty}$ as
$\frac{t^{\infty}}{\infty}\equiv\lim_{p\to\infty}\frac{t^{p}}{p}=\begin{cases}0,&0\leqslant t\leqslant 1,\\[4.30554pt] \infty,&t>1.\end{cases}$
Then $\Phi$ is a Young function,
$L^{\Phi}(\mathbf{R}^{d})$ $=L^{p_{1}}(\mathbf{R}^{d})+L^{p_{2}}(\mathbf{R}^{d}),$ $\quad p_{1}$ $\leqslant p_{2},$
and
$L^{\Phi}(\mathbf{R}^{d})$ $=L^{p_{1}}(\mathbf{R}^{d})\cap L^{p_{2}}(\mathbf{R}^{d}),$ $\quad p_{2}$ $\leqslant p_{1}.$
Example 1.9.
The previous remark shows that sums and intersections of Lebesgue spaces are special cases of Orlicz spaces. Here we list some other choices of Young functions which give rise to Orlicz spaces, where not all of them be described by Lebesgue spaces.
  • Let
    $\Phi(t)=\begin{cases}\tan t,&0\leq t<\frac{\pi}{2},\\[4.30554pt] \infty,&t\geq\frac{\pi}{2}.\end{cases}$
    It follows that $L^{\Phi}=L^{1}\cap L^{\infty}$.
  • Let
    $\Phi(t)=\begin{cases}0,&t=0,\\[4.30554pt] -\frac{t}{\ln t},&0<t<1,\\[4.30554pt] \infty,&t\geq 1,\end{cases}$
    Then the conjugate Young function is given by
    $\Phi^{*}(t)=\left(t+\frac{1}{2}-\sqrt{\frac{1}{4}+t}\,\right)e^{-\frac{1}{t}(\frac{1}{2}+\sqrt{\frac{1}{4}+t}\,)},$
    when $t\geq 0$ is near origin.
  • If $\Phi(t)=t\ln{(1+t)}$, then $\Phi^{*}(t)\asymp\cosh{(t)}-1$.
  • If $\Phi(t)=\cosh{(t)}-1$, then $\Phi^{*}(t)\asymp t\ln{(1+t)}$.
We observe that each one of these Young functions gives rise to different Orlicz spaces.
We refer to [39, 43, 34] for more facts about Orlicz spaces.
1.4. Orlicz modulation spaces
Let $\mathcal{M}(\mathbf{R}^{d})$ be the set of all (complex-valued) Lebesgue measurable functions on $\mathbf{R}^{d}$. For any $p,q\in[1,\infty]$ and $\omega\in\mathscr{P}_{E}(\mathbf{R}^{2d})$, the norm $\|\,\cdot\,\|_{M^{p,q}_{(\omega)}}$ on $\mathcal{M}(\mathbf{R}^{2d})$ is given by
$\|F\|_{L^{p,q}_{(\omega)}}\equiv\|H_{F,\omega,p}\|_{L^{q}},\qquad H_{F,\omega,p}(\xi)\equiv\|F(\,\cdot\,,\xi)\cdot\omega(\,\cdot\,,\xi)\|_{L^{q}},\qquad F\in\mathcal{M}(\mathbf{R}^{2d}).$
The definition of classical and Orlicz modulation spaces are given in the following. (See also [21, 23] for first definition of classical and more general classes of modulation spaces.)
Definition 1.10.
Let $f\in\Sigma_{1}(\mathbf{R}^{d})$, $\phi(x)=\pi^{-\frac{d}{4}}e^{-\frac{|x|^{2}}{2}},\ x\in\mathbf{R}^{d}$, $p,q\in[1,\infty]$, $\omega\in\mathscr{P}_{E}(\mathbf{R}^{2d})$, and let $\Phi$ and $\Psi$ be Young functions.
  1. (1)
    The modulation space $M^{p,q}_{(\omega)}(\mathbf{R}^{d})$ consists of all $f\in\Sigma_{1}^{\prime}(\mathbf{R}^{d})$ such that
    $\|f\|_{M^{p,q}_{(\omega)}}\equiv\|V_{\phi}\|_{L^{p,q}_{(\omega)}},$ (1.16)
    is finite. The topology of $M^{p,q}_{(\omega)}(\mathbf{R}^{d})$ is given by the norm (1.16).
  2. (2)
    The Orlicz modulation space $M^{\Phi}_{(\omega)}(\mathbf{R}^{d})$ is the set of all $f\in\mathcal{S}_{1/2}^{\prime}(\mathbf{R}^{d})$ such that
    $\|f\|_{M^{\Phi}_{(\omega)}}\equiv\|V_{\phi}f\|_{L^{\Phi}_{(\omega)}}$ (1.17)
    is finite. The topology of $M^{\Phi}_{(\omega)}(\mathbf{R}^{d})$ is given by the norm in (1.17).
Beside these well-known families of modulation spaces, we shall also consider the modulation space $M^{\sharp,q}_{(\omega)}(\mathbf{R}^{d})$, which consists of all $f\in M^{\infty,q}_{(\omega)}(\mathbf{R}^{d})$ such that
$\lim_{R\to\infty}\left(\left\|\sup_{|x|\geq R}|V_{\phi}f(x,\,\cdot\,)\omega(x,\,\cdot\,)|\right\|_{L^{q}}=0,\right)$
when $q$, $\phi$ and $\omega$ are the same as in Definition 1.10. We notice that $M^{\sharp,1}_{(\omega)}(\mathbf{R}^{d})$ is a central modulation space in [26], and that several invariance properties are deduced in [38]. For example it is here shown that the window function $\phi$ can be any element in a suitable weighted $M^{1,1}$ class. Furthermore, in [38] the following result is obtained.
Lemma 1.11.
Let $\omega\in\mathscr{P}_{E}(\mathbf{R}^{2d})$ and $q\in[1,\infty)$. Then $M^{\sharp,q}_{(\omega)}(\mathbf{R}^{d})$ is the completion of $\Sigma_{1}(\mathbf{R}^{d})$ under the norm $\|\,\cdot\,\|_{M^{\sharp,q}_{(\omega)}}$.
For convenience we set $M^{p,p}_{(\omega)}=M^{p}_{(\omega)}$. We also set
$M^{\Phi}_{(\omega)}=M^{\Phi},\quad M^{p,q}_{(\omega)}=M^{p,q}\quad\text{and}\quad M^{p}_{(\omega)}=M^{p}\quad\text{when $\omega=1$ everywhere}.$
Evidently, in Definition 1.10, we may use the transform $T_{\phi}$ in (1.5) instead of $V_{\phi}$
Let $\Phi$ be Young functions, and let $\Phi_{[p]}$ be the same as in Remark 1.8 and $\omega\in\mathscr{P}_{E}(\mathbf{R}^{2d})$. Then evidently
$M^{p}_{(\omega)}(\mathbf{R}^{d})$ $=M^{\Phi}_{(\omega)}(\mathbf{R}^{d})$ when $\quad\Phi$ $=\Phi_{[p]}.$ (1.18)
Next we explain some basic properties of Orlicz modulation spaces. The following proposition shows that Orlicz modulation spaces are completely determined by the behavior of the Young functions near origin. We refer to [63, Proposition 5.11] for the proof.
Proposition 1.12.
Let $\Phi_{j}$ be Young functions and $\omega\in\mathscr{P}_{E}(\mathbf{R}^{d})$. Then the following conditions are equivalent:
  1. (1)
    $M^{\Phi_{1}}_{(\omega)}(\mathbf{R}^{d})\subseteq M^{\Phi_{2}}_{(\omega)}(\mathbf{R}^{d})$;
  2. (2)
    for some $t_{0}>0$ it holds $\Phi_{2}(t)\lesssim\Phi_{1}(t)$ when $t\in[0,t_{0}]$.
The next two proposition show some other convenient properties concerning norm invariance and duality for Orlicz modulation spaces. We refer to Section 4 in [24] for their proofs. For an exposition with Orlicz spaces in focus, see [63]. In the unweighted case, some of the properties also follows from [50]. In the first proposition on norm invariance, we also remark that these properties hold for a significantly broader family of modulation spaces which also includes more general quasi-Banach spaces (see [61]). For classical modulation spaces, the results can be found in Chapters 11 and 12 in [28].
Proposition 1.13.
Let $\Phi$ be a Young function, $\omega,v\in\mathscr{P}_{E}(\mathbf{R}^{2d})$ be such that $\omega$ is $v$-moderate, $\phi\in M^{1}_{(v)}(\mathbf{R}^{d})\setminus 0$, and let $f\in\Sigma_{1}^{\prime}(\mathbf{R}^{d})$. Then $f\in M^{\Phi}_{(\omega)}(\mathbf{R}^{d})$, if and only if
$\|f\|\equiv\|V_{\phi}f\cdot\omega\|_{L^{\Phi}}$ (1.19)
is finite. Furthermore, $\|\,\cdot\,\|$ defines a norm on $M^{\Phi}_{(\omega)}(\mathbf{R}^{d})$ which is equivalent to $\|\,\cdot\,\|_{M^{\Phi}_{(\omega)}}$.
Proposition 1.14.
Let $\Phi$ be a Young function, and let $\omega\in\mathscr{P}_{E}(\mathbf{R}^{2d})$. Then the following is true:
  1. (1)
    the sesqui-linear form $(\,\cdot\,,\,\cdot\,)_{L^{2}}$ on $\Sigma_{1}(\mathbf{R}^{d})$ extends to a continuous map from
    $M^{\Phi}_{(\omega)}(\mathbf{R}^{d})\times M^{\Phi^{*}}_{(1/\omega)}(\mathbf{R}^{d})$
    to $\mathbf{C}$. This extension is unique when $\Phi$ and $\Psi$ fulfill a local $\Delta_{2}$-condition. If $\|f\|=\sup|{(f,g)_{L^{2}}}|$, where the supremum is taken over all $b\in M^{\Phi^{*}}_{(1/\omega)}(\mathbf{R}^{d})$ such that $\|b\|_{M^{\Phi^{*}}_{(1/\omega)}}\leq 1$, then $\|\cdot\|$ and $\|\cdot\|_{M^{\Phi}_{(\omega)}}$ are equivalent norms;
  2. (2)
    if $\Phi$ and $\Psi$ fulfill a local $\Delta_{2}$-condition, then $\Sigma_{1}(\mathbf{R}^{d})$ is dense in $M^{\Phi}_{(\omega)}(\mathbf{R}^{d})$, and the dual space of $M^{\Phi}_{(\omega)}(\mathbf{R}^{d})$ can be identified with $M^{\Phi^{*}}_{(1/\omega)}(\mathbf{R}^{d})$, through the form $(\,\cdot\,,\,\cdot\,)_{L^{2}}$. Moreover, $\Sigma_{1}(\mathbf{R}^{d})$ is weakly dense in $M^{\Phi^{*}}_{(\omega)}(\mathbf{R}^{d})$.
Remark 1.15.
We notice that the weight classes play important roles for the sizes of modulation spaces. More precisely, let $\Phi$ be a Young function and $\omega$ be a weight on $\mathbf{R}^{2d}$. Then
$\Sigma_{s}(\mathbf{R}^{d})$ $\hookrightarrow M^{\Phi}_{(\omega)}(\mathbf{R}^{d})\hookrightarrow\Sigma_{s}^{\prime}(\mathbf{R}^{d}),$ when $\quad\omega$ $\in\mathscr{P}_{s}(\mathbf{R}^{2d}),$ $\quad s$ $\in(0,1],$
$\mathcal{S}_{s}(\mathbf{R}^{d})$ $\hookrightarrow M^{\Phi}_{(\omega)}(\mathbf{R}^{d})\hookrightarrow\mathcal{S}_{s}^{\prime}(\mathbf{R}^{d}),$ when $\quad\omega$ $\in\mathscr{P}_{0,s}(\mathbf{R}^{2d}),$ $\quad s$ $\in(0,1],$
and
$\mathscr{S}(\mathbf{R}^{d})$ $\hookrightarrow M^{\Phi}_{(\omega)}(\mathbf{R}^{d})\hookrightarrow\mathscr{S}^{\prime}(\mathbf{R}^{d}),$ when $\quad\omega$ $\in\mathscr{P}(\mathbf{R}^{2d}).$
These embeddings are narrow in the sense
$\bigcap_{\omega\in\mathscr{P}_{s}}M^{\Phi}_{(\omega)}(\mathbf{R}^{d})$ $=\Sigma_{s}(\mathbf{R}^{d}),$ $\qquad\bigcup_{\omega\in\mathscr{P}_{s}}M^{\Phi}_{(\omega)}(\mathbf{R}^{d})$ $=\Sigma_{s}^{\prime}(\mathbf{R}^{d}),$ $\quad s$ $\in(0,1],$
$\bigcap_{\omega\in\mathscr{P}_{0,s}}M^{\Phi}_{(\omega)}(\mathbf{R}^{d})$ $=\mathcal{S}_{s}(\mathbf{R}^{d}),$ $\qquad\bigcup_{\omega\in\mathscr{P}_{0,s}}M^{\Phi}_{(\omega)}(\mathbf{R}^{d})$ $=\mathcal{S}_{s}^{\prime}(\mathbf{R}^{d}),$ $\quad s$ $\in(0,1],$
and
$\bigcap_{\omega\in\mathscr{P}}M^{\Phi}_{(\omega)}(\mathbf{R}^{d})$ $=\mathscr{S}(\mathbf{R}^{d}),$ $\qquad\bigcup_{\omega\in\mathscr{P}}M^{\Phi}_{(\omega)}(\mathbf{R}^{d})$ $=\mathscr{S}^{\prime}(\mathbf{R}^{d}).$
(See e. g. [63, 61].)
Example 1.16.
Let $\phi\in\mathscr{S}(\mathbf{R}^{d})\setminus 0$, and let $\Phi$ be convex on $0,\infty)$ such that
$\Phi(t)=-t^{2}\ln t,\quad\text{when}\quad t\in[0,e^{-\frac{2}{3}}].$
Then $\Phi$ is a Young function. The entropy functional
$E_{\phi}(f)\equiv-\iint_{\mathbf{R}^{2d}}|V_{\phi}f(x,\xi)|^{2}\ln|V_{\phi}f(x,\xi)|^{2}\,dxd\xi$ (1.20)
is appears when investigating kinetic energy in statistical physics and quantum physics, see e. g. [36].
It is proved in [33] that the following
  1. (1)
    The space $M^{\Phi}(\mathbf{R}^{d})$ is close to $M^{2}(\mathbf{R}^{d})$ in the sense of the continuous inclusions
    $M^{p}(\mathbf{R}^{d})\subseteq M^{\Phi}(\mathbf{R}^{d})\underset{\text{dense}}{\subseteq}M^{2}(\mathbf{R}^{d}),\qquad p<2.$
  2. (2)
    The functional $E_{\phi}$ is continuous on $M^{\Phi}(\mathbf{R}^{d})$, but fails to be continuous on $M^{2}(\mathbf{R}^{d})$.
As a consequence of (1) and (2) one has that $E_{\phi}$ is continuous on $M^{p}(\mathbf{R}^{d})$ when $p<2$, which seems not to be known before [33].
1.5. Kernel operators and Schatten-von Neumann classes
For any topological vector spaces, $V_{1}$ and $V_{2}$, we let $\mathcal{L}(V_{1},V_{2})$ be the set of all linear and continuous operators from $V_{1}$ to $V_{2}$. Suppose $V_{1}=\Sigma_{1}(\mathbf{R}^{d_{1}})$, $V_{2}=\Sigma_{1}^{\prime}(\mathbf{R}^{d_{2}})$ and $K\in\Sigma_{1}^{\prime}(\mathbf{R}^{d_{2}}\times\mathbf{R}^{d_{1}})$. Then it follows that the map $T_{K}$, defined by
$\langle T_{K}f,g\rangle=\langle K,g\otimes f\rangle,\qquad f\in\Sigma_{1}(\mathbf{R}^{d_{1}}),\ g\in\Sigma_{1}(\mathbf{R}^{d_{2}}),$
belongs to $\mathcal{L}(V_{1},V_{2})$. By the kernel theorem of Schwartz it follows that the map $K\mapsto T_{K}$, from $\Sigma_{1}^{\prime}(\mathbf{R}^{d_{2}}\times\mathbf{R}^{d_{1}})$ to $\mathcal{L}(V_{1},V_{2})$ is bijective. For convenience we let the topology of $\mathcal{L}(V_{1},V_{2})$ be inherited from the topology of $\Sigma_{1}^{\prime}(\mathbf{R}^{d_{2}}\times\mathbf{R}^{d_{1}})$.
In what follows we recall some facts on Schatten-von Neumann operators, given in [51]. Let $T$ be a linear and continuous map from the Hilbert space $\mathcal{H}_{1}$ into the Hilbert space $\mathcal{H}_{2}$, and let $j\geq 1$ be an integer. Also let $\mathscr{I}_{0,j}(\mathcal{H}_{1},\mathcal{H}_{2})$ be the set of all linear and continuous operators from $\mathcal{H}_{1}$ to $\mathcal{H}_{2}$ with rank at most j1j-1. The singular value of $T$ of order $j$ is defined by
$\sigma_{j}(T)=\sigma_{j}(T;\mathcal{H}_{1},\mathcal{H}_{2})\equiv\inf_{T_{0}\in\mathscr{I}_{0,j}}\|T-T_{0}\|_{\mathcal{H}_{1}\to\mathcal{H}_{2}},\qquad j\in\mathbf{Z}_{+}.$
Evidently $\sigma_{j}(T;\mathcal{H}_{1},\mathcal{H}_{2})$ decreases with $j$, and $\sigma_{1}(T;\mathcal{H}_{1},\mathcal{H}_{2})$ is equal to the operator norm $\|T\|_{\mathcal{H}_{1}\to\mathcal{H}_{2}}$ of $T$.
Throughout the paper, all Hilbert spaces are assumed to be separable, and observe that this is always the case for Hilbert spaces which are continuously embedded in $\mathscr{S}^{\prime}(\mathbf{R}^{d})$, in view of [44, Proposition 1.2]. However, we note that most parts of what is described here also hold when $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are allowed to be non-separable.
In the following definition we present a broad family of Schatten-von Neumann classes.
Definition 1.17.
Let $\mathcal{H}_{1}$, $\mathcal{H}_{2}$ be Hilbert spaces, $T$ be a linear operator from $\mathcal{H}_{1}$ to $\mathcal{H}_{2}$, and let $\mathcal{B}\subseteq\ell_{0}^{\prime}(\mathbf{Z}_{+})$ be a Banach space.
  1. (1)
    The $\mathcal{B}$ Schatten-von Neumann norm of $T$ is given by
    $\|T\|_{\mathscr{I}_{\mathcal{B}}}=\|T\|_{\mathscr{I}_{\mathcal{B}}(\mathcal{H}_{1},\mathcal{H}_{2})}\equiv\|\{\sigma_{j}(T;\mathcal{H}_{1},\mathcal{H}_{2})\}_{j=1}^{\infty}\|_{\mathcal{B}}.$
  2. (2)
    The $\mathcal{B}$ Schatten-von Neumann class $\mathscr{I}_{\mathcal{B}}=\mathscr{I}_{\mathcal{B}}(\mathcal{H}_{1},\mathcal{H}_{2})$ consists of all linear and continuous operators $T$ from $\mathcal{H}_{1}$ to $\mathcal{H}_{2}$ such that $\|T\|_{\mathscr{I}_{\mathcal{B}}(\mathcal{H}_{1},\mathcal{H}_{2})}$ is finite. The topology of $\mathscr{I}_{\mathcal{B}}=\mathscr{I}_{\mathcal{B}}(\mathcal{H}_{1},\mathcal{H}_{2})$ is given through the norm $\|\,\cdot\,\|_{\mathscr{I}_{\mathcal{B}}(\mathcal{H}_{1},\mathcal{H}_{2})}$.
Definition 1.18.
Let $\Phi$ be a Young function, $p\in[1,\infty]$ and let $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ be Hilbert spaces. Then let
$\mathscr{I}_{\Phi}=\mathscr{I}_{\ell^{\Phi}},\quad\mathscr{I}_{p}=\mathscr{I}_{\ell^{p}}\quad\text{and}\quad\mathscr{I}_{\sharp}=\mathscr{I}_{\ell^{\sharp}}.$
  • The space $\mathscr{I}_{\Phi}(\mathcal{H}_{1},\mathcal{H}_{2})$ is called the Orlicz Schatten-von Neumann class with respect to $\Phi$, $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$, or the $\Phi$-Schatten class.
  • The space $\mathscr{I}_{p}(\mathcal{H}_{1},\mathcal{H}_{2})$ is called the (classical) Schatten-von Neumann class with respect to $p$, $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$, or the $p$-Schatten class.
We observe that $\mathscr{I}_{p}(\mathcal{H}_{1},\mathcal{H}_{2})$ increases with $p$, and that $\mathscr{I}_{\Phi}(\mathcal{H}_{1},\mathcal{H}_{2})$ decreases with $\Phi$. In fact, for the latter conclusion, it suffices to detect the decreasing property with respect to $\Phi$ near origin, which is shown in the following proposition. The result follows from the fact that similar properties hold true for discrete Orlicz spaces. (See e. g. [3, 50].) The details are left for the reader.
Proposition 1.19.
Let $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ be Hilbert spaces, and let $\Phi_{1}$ and $\Phi_{2}$ be Young functions such that for some $T>0$ it holds
$\Phi_{2}(t)\lesssim\Phi_{1}(t),\quad\text{when}\quad t\in(0,T].$
Then $\mathscr{I}_{\Phi_{1}}(\mathcal{H}_{1},\mathcal{H}_{2})\hookrightarrow\mathscr{I}_{\Phi_{2}}(\mathcal{H}_{1},\mathcal{H}_{2})$.
We notice that
$\mathscr{I}_{1}(\mathcal{H}_{1},\mathcal{H}_{2}),\quad\mathscr{I}_{2}(\mathcal{H}_{1},\mathcal{H}_{2}),\quad\text{and}\quad\mathscr{I}_{\infty}(\mathcal{H}_{1},\mathcal{H}_{2}),$
are the spaces of trace-class, Hilbert-Schmidt, and linear and continuous operators from $\mathcal{H}_{1}$ to $\mathcal{H}_{2}$, respectively, also in norms.
By straight-forward application of the spectral theorem, it follows that the definition of $\mathscr{I}_{p}(\mathcal{H})$, $\mathscr{I}_{\Phi}(\mathcal{H})$ and their norms can be reached with other approaches, indicated in the following.
Proposition 1.20.
Let $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ be Hilbert spaces, and $\Phi$ be a Young function. Then $\mathscr{I}_{\Phi}(\mathcal{H}_{1},\mathcal{H}_{2})$ is a Banach, and
$\|T\|_{\mathscr{I}_{\Phi}}=\sup\|\{(Tf_{j},g_{j})_{\mathcal{H}_{2}}\}_{j=1}^{\infty}\|_{\ell^{\Phi}(\mathbf{Z}_{+})},\quad T\in\mathscr{I}_{\infty}(\mathcal{H}_{1},\mathcal{H}_{2}).$ (1.21)
Here the supremum is taken over all orthonormal sequences $\{f_{j}\}_{j=1}^{\infty}\in\operatorname{ON}(\mathcal{H}_{1})$ and $\{g_{j}\}_{j=1}^{\infty}\in\operatorname{ON}(\mathcal{H}_{2})$.
Since our investigations especially concerns Schatten-von Neumann properties for operators acting on Hilbert modulation spaces, it is convenient for us to set
$\mathscr{I}_{\Phi}(\omega_{1},\omega_{2})=\mathscr{I}_{\Phi}(M^{2}_{(\omega_{1})}(\mathbf{R}^{d_{1}}),M^{2}_{(\omega_{2})}(\mathbf{R}^{d_{2}}))$
and
$\mathscr{I}_{p}(\omega_{1},\omega_{2})=\mathscr{I}_{p}(M^{2}_{(\omega_{1})}(\mathbf{R}^{d_{1}}),M^{2}_{(\omega_{2})}(\mathbf{R}^{d_{2}})).$
2. Continuity for Fourier integral operators
In this section we obtain continuity for Fourier integral operators when acting on Orlicz modulation spaces. In the first part (Subsection 2.1) we deduce continuity properties for operators with kernels belonging to Orlicz modulation spaces. Thereafter we consider Fourier integral operators with amplitudes belonging to $M^{\infty,1}_{(\omega)}$ and the phase function $\varphi$ should satisfy (0.2) and $\varphi^{\prime\prime}\in M^{\infty,1}_{(v)}$. Here $\omega$ is $v$-moderate. We show that such Fourier integral operators are continuous from $M^{\Phi}_{(\omega_{1})}$ to $M^{\Phi}_{(\omega_{2})}$, provided the weight functions $\omega$, $\omega_{1}$ and $\omega_{2}$ obey suitable estimate conditions.
Thereafter we consider Fourier integral operators with amplitudes belonging to the Orlicz modulation space $M^{\Phi}_{(\omega)}$, and where the condition (0.2) is replaced by other ones. We show that such operators are continuous from $M^{\Phi^{*}}_{(\omega_{1})}$ to $M^{\Phi}_{(\omega_{2})}$.
2.1. Mapping properties for certain kernel operators on Orlicz modulation spaces
The next result concerns mapping properties for operators with kernels in Orlicz modulation spaces.
Proposition 2.1.
Let $\Phi$ be a Young function such that at least one of $\Phi$ and $\Phi^{*}$ satisfies the $\Delta_{2}$-condition, let $\omega\in\mathscr{P}_{E}(\mathbf{R}^{2d})$ and let $\omega_{j}\in\mathscr{P}_{E}(\mathbf{R}^{d_{j}})$, $j=1,2$, be such that
$\frac{\omega_{2}(x_{2},\xi_{2})}{\omega_{1}(x_{1},-\xi_{1})}\lesssim\omega(x_{1},x_{2},\xi_{1},\xi_{2}),\qquad x_{j},\xi_{j}\in\mathbf{R}^{d_{j}}.$
Also let $K\in M^{\Phi}_{(\omega)}(\mathbf{R}^{d_{2}+d_{1}})$. Then $T_{K}$ from $\Sigma_{1}(\mathbf{R}^{d_{1}})$ to $\Sigma_{1}^{\prime}(\mathbf{R}^{d_{2}})$ extends uniquely to a continuous operator from $M^{\Phi^{*}}_{(\omega_{1})}(\mathbf{R}^{d_{1}})$ to $M^{\Phi}_{(\omega_{2})}(\mathbf{R}^{d_{2}})$,
$\|T_{K}f\|_{M^{\Phi}_{(\omega_{2})}}$ $\leq C\|K\|_{M^{\Phi}_{(\omega)}}\|f\|_{M^{\Phi^{*}}_{(\omega_{1})}},$ (2.1)
$K$ $\in M^{\Phi}_{(\omega)}(\mathbf{R}^{d_{2}+d_{1}}),\quad f\in M^{\Phi^{*}}_{(\omega_{1})}(\mathbf{R}^{d_{1}}).$
The result follows by standard arguments in operator theory. In order to assist the reader we here present the arguments.
Proof.
First suppose $\Phi$ satisfies the $\Delta_{2}$- condition. Let $\phi_{j}\in\Sigma_{1}(\mathbf{R}^{d_{j}})\setminus 0$ and let $\phi(x,y)=\phi_{2}(x)\overline{\phi_{1}(y)}$. Since $\mathscr{S}(\mathbf{R}^{d_{2}+d_{1}})$ is dense in $M^{\Phi}_{(\omega)}(\mathbf{R}^{d_{2}+d_{1}})$, it suffices to prove (2.1) when $K\in\Sigma_{1}(\mathbf{R}^{d_{2}+d_{1}})$. Let
$\mathcal{K}(x,\xi,y,\eta)=V_{\phi}K(x,y,\xi,-\eta)\omega(x,y,\xi,-\eta)\quad\text{and}\quad F=V_{\phi_{1}}f\omega_{1}.$
Then
$|V_{\phi_{2}}(T_{K}f)(x,\xi)\omega_{2}(x,\xi)|\lesssim|(\mathcal{K}(x,\xi,\,\cdot\,),F))_{L^{2}(\mathbf{R}^{2d_{2}})}|\lesssim\|\mathcal{K}(x,\xi,\,\cdot\,)\|_{L^{\Phi}}\|F\|_{L^{\Phi^{*}}}.$
This gives
$\|T_{K}f\|_{M^{\Phi}_{(\omega_{2})}}\lesssim\|\mathcal{K}\|_{L^{\Phi}}\|F\|_{L^{\Phi^{*}}}\asymp\|K\|_{M^{\Phi}_{(\omega)}}\|f\|_{M^{\Phi^{*}}_{(\omega_{1})}}.$
The case when $\Phi^{*}$ satisfies the $\Delta_{2}$-condition follows by straight-forward modification of the arguments, giving the estimate
$|(T_{K}f,g)_{L^{2}(\mathbf{R}^{d_{2}})}|\lesssim\|K\|_{M^{\Phi}_{(\omega)}}\|f\|_{M^{\Phi^{*}}_{(\omega_{1})}}\|g\|_{M^{\Phi^{*}}_{(1/\omega_{2})}},$
when $f\in\Sigma_{1}(\mathbf{R}^{d_{1}})$ and $g\in\Sigma_{1}(\mathbf{R}^{d_{2}})$, and then using the fact that $\mathscr{S}$ is dense in $M^{\Phi^{*}}_{(\omega_{1})}$ and in $M^{\Phi^{*}}_{(1/\omega_{2})}$. The details are left for the reader. ∎
2.2. Continuity for Fourier integral operators, when acting on modulation spaces
In order to explain our results on Fourier integral operators, we first give the conditions on involved weight functions. We usually assume that they satisfy
$\frac{\omega_{2}(x,\xi)}{\omega_{1}(y,-\eta)}$ $\leq C_{1}\omega_{0}(x,y,\xi,\eta)$ (2.2)
$\leq C_{2}\omega(X,\xi-\varphi^{\prime}_{x}(X),\eta-\varphi^{\prime}_{y}(X),-\varphi^{\prime}_{\zeta}(X)),$
$\omega(X,\xi_{1}+$ $\xi_{2},\eta_{1}+\eta_{2},z_{1}+z_{2})\leq C\omega(X,\xi_{1},\eta_{1},z_{1})v_{0}(\xi_{2},\eta_{2},z_{2}),$
$\omega$ $\in\mathscr{P}_{s}(\mathbf{R}^{2(2N+m)}),\quad\omega_{0}\in\mathscr{P}_{s}(\mathbf{R}^{2(2N)}),$
$\omega_{j}$ $\in\mathscr{P}_{s}(\mathbf{R}^{2d_{j}}),\quad N=d_{1}+d_{2},\ j=1,2,$
$v_{0}$ $\in\mathscr{P}_{s}(\mathbf{R}^{2N+m}),\quad\sup_{t\in[0,1]}\|v_{0}(t\,\cdot\,)/v_{0}\|_{L^{\infty}}<\infty,$
$v(X,\xi,\eta,z)$ $=v_{0}(\xi,\eta,z),\ X=(x,y,\zeta)\in\mathbf{R}^{2N+m},$
and that the phase function should satisfy
$\varphi\in C(\mathbf{R}^{2d+m})\quad\text{and}\quad\varphi^{(\alpha)}\in M^{\infty,1}_{(v)}(\mathbf{R}^{2d+m}),\ |\alpha|=2.$ (2.3)
The first result shows that Fourier integral operators with amplitudes in $M^{\infty,1}_{(\omega)}(\mathbf{R}^{2d+N})$ are well-defined as continuous mappings from $\Sigma_{1}(\mathbf{R}^{d})$ to $\Sigma_{1}^{\prime}(\mathbf{R}^{d})$. For polynomial weights, the result is essentially the same as [60, Theorem 2.1]. For general weights, the result follows by similar arguments as for [60, Theorem 2.1]. The proof is therefore omitted.
Theorem 2.2.
Let $s>1$, $d_{1}=d_{2}=d$, $\varphi\in C(\mathbf{R}^{2d+m})$, and $\omega,v\in\mathscr{P}_{s}(\mathbf{R}^{2(2d+m)})$ satisfy (0.2), (2.2) and (2.3). Then the map $a\mapsto\operatorname{Op}_{\varphi}(a)$ from $\Sigma_{1}(\mathbf{R}^{2d+m})$ to $\mathcal{L}(\Sigma_{1}(\mathbf{R}^{d}),\Sigma_{1}^{\prime}(\mathbf{R}^{d}))$ extends uniquely to a continuous map from $M^{\infty,1}_{(\omega)}(\mathbf{R}^{2d+m})$ to $\mathcal{L}(\Sigma_{1}(\mathbf{R}^{d}),\Sigma_{1}^{\prime}(\mathbf{R}^{d}))$.
The next two theorems assert continuity and compactness for Fourier integral operators in Theorem 2.2, when they are acting on Orlicz modulation spaces.
Theorem 2.3.
Let $s>1$, $d_{1}=d_{2}=d$, $\Phi$ be a Young function such that $1<q_{\Phi}\leq p_{\Phi}<\infty$, $\omega_{1},\omega_{2}\in\mathscr{P}_{s}(\mathbf{R}^{2d})$, and let $\varphi$, $\omega$, $v$ be the same as in Theorem 2.2. Also let $a\in M^{\infty,1}_{(\omega)}(\mathbf{R}^{2d+m})$. Then $\operatorname{Op}_{\varphi}(a)$ from $\Sigma_{1}(\mathbf{R}^{d})$ to $\Sigma_{1}^{\prime}(\mathbf{R}^{d})$ extends uniquely to a continuous operator from $M^{\Phi}_{(\omega_{1})}(\mathbf{R}^{d})$ to $M^{\Phi}_{(\omega_{2})}(\mathbf{R}^{d})$. Moreover, for some constant $C$ it holds
[…]
Theorem 2.4.
Let $s>1$, $d_{1}=d_{2}=d$, $\Phi$ be a Young function such that $1<q_{\Phi}\leq p_{\Phi}<\infty$, $\omega_{1},\omega_{2}\in\mathscr{P}_{s}(\mathbf{R}^{2d})$, and let $\varphi$, $\omega$, $v$ be the same as in Theorem 2.2. Also let $a\in M^{\sharp,1}_{(\omega)}(\mathbf{R}^{2d+m})$. Then $\operatorname{Op}_{\varphi}(a)$ from $M^{\Phi}_{(\omega_{1})}(\mathbf{R}^{d})$ to $M^{\Phi}_{(\omega_{2})}(\mathbf{R}^{d})$ is compact.
Theorem 2.4 will follow by combining Theorem 2.3, with certain density arguments and the following proposition.
Proposition 2.5.
Let $s>1$, $\Phi$, $\varphi$, $\omega$, $v$, $\omega_{1}$, $\omega_{2}$ be the same as in Theorem 2.3, and let $a\in\Sigma_{1}(\mathbf{R}^{2d+m})$. Then $\operatorname{Op}_{\varphi}(a)$ from $M^{\Phi}_{(\omega_{1})}(\mathbf{R}^{d})$ to $M^{\Phi}_{(\omega_{2})}(\mathbf{R}^{d})$ is compact.
Before proving Theorems 2.3, 2.4, and Proposition 2.5, we present some further continuity properties for Fourier integral operators. The amplitudes should fulfill norm estimates, with norm $\|\,\cdot\,\|$ either defined by
$\|a\|$ $=\left\|\int_{\mathbf{R}^{m}}H_{a,\omega}(\,\cdot\,,\zeta)\,d\zeta\right\|_{L^{\Phi}},$ (2.5)
$H_{a,\omega}(X)$ $=\sup_{z\in\mathbf{R}^{m}}\left(\|V_{\phi}a(X,\,\cdot\,,z)\omega(X,\,\cdot\,,z)\|_{L^{\Phi}}\right),\quad X=(x,y,\zeta),$
or
$\|a\|$ $=\left\|\int_{\mathbf{R}^{m}}H_{a,\omega}(\,\cdot\,,z)\,dz\right\|_{L^{\Phi}},$ (2.6)
$H_{a,\omega}(x,y,z)$ $=\sup_{\zeta\in\mathbf{R}^{m}}\left(\|V_{\phi}a(X,\,\cdot\,,z)\omega(X,\,\cdot\,,z)\|_{L^{\Phi}}\right),\quad X=(x,y,\zeta),$
or by
$\|a\|$ $=\int_{\mathbf{R}^{m}}\|H_{a,\omega}(\,\cdot\,,z)\|_{L^{\Phi}}\,dz,$ (2.7)
$H_{a,\omega}(x,y,\xi,\eta,z)$ $=\sup_{\zeta\in\mathbf{R}^{m}}\left(|V_{\phi}a(X,\xi,\eta,z)\omega(X,\xi,\eta,z)|\right),\quad X=(x,y,\zeta).$
We search estimates of the form
$\|K_{a,\varphi}\|_{M^{\Phi}_{(\omega_{0})}}\leq C{d\hskip-0.80002pt\bar{}\hskip 1.99997pt}^{-1}\exp\left(\|\varphi^{\prime\prime}\|_{M^{\infty,1}_{(v)}}\right)\|a\|,$ (2.8)
and
$\|\operatorname{Op}_{\varphi}(a)\|_{M^{\Phi^{*}}_{(\omega_{1})}\to M^{\Phi}_{(\omega_{2})}}\leq C{d\hskip-0.80002pt\bar{}\hskip 1.99997pt}^{-1}\exp\left(\|\varphi^{\prime\prime}\|_{M^{\infty,1}_{(v)}}\right)\|a\|.$ (2.9)
Theorem 2.6.
Let $s>1$, $N=d_{2}+d_{1}$, $\Phi$ be a Young function such that $1<q_{\Phi}\leq p_{\Phi}<\infty$, $q_{\Phi}=p_{\Phi}=1$ or $q_{\Phi}=p_{\Phi}=\infty$, let $\phi\in\Sigma_{s}(\mathbf{R}^{N+m})\setminus 0$, $\omega$, $\omega_{j}$, $j=0,1,2$, $v$ and $\varphi$ be as in (2.2) and (2.3). Also let $a\in\Sigma_{s}^{\prime}(\mathbf{R}^{N+m})$ be such that one of the following conditions holds:
  • (i)
    $\|a\|$ in (2.5) is finite;
  • (ii)
    $\|a\|$ in (2.6) is finite and that $|\det(\varphi^{\prime\prime}_{\zeta,\zeta})|\geq d\hskip-0.80002pt\bar{}\hskip 1.99997pt$ for some $d\hskip-0.80002pt\bar{}\hskip 1.99997pt>0$.
Then the following is true:
  1. (1)
    the kernel $K_{a,\varphi}$ of $\operatorname{Op}_{\varphi}(a)$ belongs to $M^{\Phi}_{(\omega_{0})}(\mathbf{R}^{d})$, and (2.8) holds for some constant $C$ which is independent of $a\in\Sigma_{s}^{\prime}(\mathbf{R}^{N+m})$ and $\varphi\in C(\mathbf{R}^{N+m})$;
  2. (2)
    the definition of $\operatorname{Op}_{\varphi}(a)$ extends uniquely to a continuous operator from $M^{\Phi^{*}}_{(\omega_{1})}(\mathbf{R}^{d_{1}})$ to $M^{\Phi}_{(\omega_{2})}(\mathbf{R}^{d_{2}})$, and (2.9) holds.
For the proofs of these results we need the following lemma. Here we formulate the action of $\operatorname{Op}_{\varphi}(a)$ as
$(\operatorname{Op}_{\varphi}(a)f_{1},f_{2})$ $=\iiint_{\mathbf{R}^{4N+m}}\mathcal{K}_{a,\varphi}(X,\xi,\eta)F_{1}(y,\eta)\overline{F_{2}(x,\xi)}e^{-i(\langle x,\xi\rangle+\langle y,\eta\rangle)}\,dXd\xi d\eta,$ (2.10)
with
$F_{1}(y,\eta)$ $=V_{\phi_{1}}f_{1}(y,-\eta)\omega_{1}(y,-\eta),\ F_{2}(x,\xi)=V_{\phi_{2}}f_{2}(x,\xi)/\omega_{2}(x,\xi)$ (2.11)
Lemma 2.7.
Let $s>1$, $\Phi$ and $\Psi$ be Young functions, $N=d_{1}+d_{2}$, $\omega,v\in\mathscr{P}_{s}(\mathbf{R}^{2(N+m)})$ be as in (2.2), $f_{j}\in\Sigma_{1}(\mathbf{R}^{d_{j}})$ and $\phi\in C_{0}^{\infty}(\mathbf{R}^{N+m})\setminus 0$. Also let $a\in M^{\Phi,\Psi}_{(\omega)}(\mathbf{R}^{N+m})$ or $a\in M^{\Phi}_{(\omega)}(\mathbf{R}^{N+m})$, and $f_{j}\in\mathscr{S}(\mathbf{R}^{d_{j}})$, $j=1,2$. Then (2.10) and (2.11) hold for some $\mathcal{K}_{a,\varphi}$, which satisfies
$|\mathcal{K}_{a,\varphi}(X,\xi,\eta)|\leq(G*|V_{\phi}a(X,\,\cdot\,)\omega(X,\,\cdot\,)|)(\xi-\varphi^{\prime}_{x}(X),\eta-\varphi^{\prime}_{y}(X),-\varphi^{\prime}_{\zeta}(X)),$
for some non-negative $G\in L^{1}(\mathbf{R}^{N+m})$ which satisfies
[…]
Proof.
Let
$\Theta=(\xi-\varphi^{\prime}_{x}(X),\eta-\varphi^{\prime}_{y}(X),-\varphi^{\prime}_{\zeta}(X)),$
and choose $\phi_{j}\in\mathcal{D}_{0,s}(\mathbf{R}^{d_{j}})$ such that
$0$ $\leq\phi_{j},\phi,$ $\quad\int_{\mathbf{R}^{d_{j}}}\phi_{j}(x_{j})\,dx_{j}$ $=\int_{\mathbf{R}^{N+m}}\phi(X)\phi_{1}(y)\phi_{2}(x)\,dX=1,$ (2.13)
$j$ $=1,2,$ $\qquad X$ $=(x,y,\zeta)\in\mathbf{R}^{N+m}.$
By (2.5), (2.15), Lemma 2.2, Lemma 2.3 in [60], and straight-forward computations, it follows that (2.10) holds with $\mathcal{K}_{a,\varphi}^{0}$,
$F_{0,1}(y,\eta)=V_{\phi_{1}}f_{1}(y,-\eta)\quad\text{and}\quad F_{0,2}(x,\xi)=V_{\phi_{2}}f_{2}(x,\xi)$
in place of $\mathcal{K}_{a,\varphi}$, $F_{1}$ and $F_{2}$, where $\mathcal{K}_{a,\varphi}^{0}$ satisfies
$|\mathcal{K}_{a,\varphi}^{0}(X,\xi,\eta)|\leq(G_{0}*|V_{\phi}a(X,\,\cdot\,)|)(\Theta),$
with $G_{0}\in L^{1}_{(v_{0})}(\mathbf{R}^{d+m})$ satisfying
[…]
The details are left for the reader. Then (2.10) holds if
$\mathcal{K}_{a,\varphi}(X,\xi,\eta)=\mathcal{K}_{a,\varphi}^{0}(X,\xi,\eta)\omega_{2}(x,\xi)/\omega_{1}(y,-\eta).$
A combination of the latter relationships with (2.2) gives
$|\mathcal{K}_{a,\varphi}(X,\xi,\eta)|$ $\leq(G_{0}*|V_{\phi}a(X,\,\cdot\,)|)(\Theta)\omega_{2}(x,\xi)/\omega_{1}(y,-\eta)$
$\lesssim(G_{0}*|V_{\phi}a(X,\,\cdot\,)|)(\Theta)\omega(X,\Theta)$
$\lesssim(G*|V_{\phi}a(X,\,\cdot\,)\omega(X,\,\cdot\,)|)(\Theta),$
where $G=G_{0}v_{0}$ satisfies (2.12). This gives the result. ∎
Remark 2.8.
Let $a$, $\phi$, $\phi_{j}$, $\varphi$ and $\Theta$ be the same as in Lemma 2.7 and its proof, let $v$ and $v_{0}$ be as in (2.2) and let
$v_{1}(x,y,\zeta,\Xi)=v_{0}(x,y,\zeta).$
Then [60, Lemma 2.2] shows that
$\mathcal{K}_{a,\varphi}(X,\xi,\eta)\cdot\frac{\omega_{1}(y,-\eta)}{\omega_{2}(x,\xi)}$ $=(2\pi)^{\frac{m}{2}}((\mathscr{F}(e^{i\varphi_{2,X}}\phi))*(T_{\phi}a))(\Theta)$ (2.14)
$\varphi_{2,X}$ $=\psi(X)\int_{0}^{1}(1-t)\langle\varphi^{\prime\prime}(X+tY)Y,Y\rangle\,dt,$ (2.15)
$\psi$ $\in C_{0}^{\infty}(\mathbf{R}^{d+m}),\quad\psi=1\ \text{on}\ \operatorname{supp}\phi.$ (2.16)
Here $T_{\phi}$ is given by (1.5). The function $G_{0}$ in the proof of Lemma 2.7 can then be chosen as
$G_{0}=|\mathscr{F}(e^{i\varphi_{2,X}}\phi)|$
By [60, Lemma 2.3] one has
$|\mathscr{F}(e^{i\varphi_{2,X}})|\leq(2\pi)^{\frac{1}{2}(d+m)}\delta_{0}+\hat{b}$,
where $b$ satisfies
[…]
It follows that
[…]
Since
$\widehat{b}\in M^{1}_{(v_{1})}(\mathbf{R}^{d+m})\subseteq L^{1}_{(v_{0})}(\mathbf{R}^{d+m})\quad\text{and}\quad\widehat{\phi}\in\mathscr{S}(\mathbf{R}^{d+m})\subseteq L^{1}_{(v_{0})}(\mathbf{R}^{d+m}),$
it follows that $G_{0}\in L^{1}_{(v_{0})}(\mathbf{R}^{d+m})$, as it was used in the proof of Lemma 2.7.
Remark 2.9.
Let $\phi$ and $\phi_{j}$ be as in Lemma 2.7, and choose $\psi_{j}\in\mathcal{D}_{0,s}(\mathbf{R}^{d_{j}})$ such that $(\phi_{j},\psi_{j})=1$, $j=1,2$. Also let $\psi=\psi_{2}\otimes\psi_{1}$. Then it follows from (2.10) and an application of Fourier’s inversion formula that
$V_{\psi}K_{a,\varphi}(x,y,\xi,\eta)\cdot\frac{\omega_{2}(x,\xi)}{\omega_{1}(y,-\eta)}\\[4.30554pt] =Ce^{-i(\langle x,\xi\rangle+\langle y,\eta\rangle)}\int_{\mathbf{R}^{m}}\mathcal{K}_{a,\varphi}(X,\xi,\eta)\,d\zeta,\quad X=(x,y,\zeta).$ (2.17)
Proof of Theorem 2.3.
We shall follow the proof of [60, Theorem 2.1] when proving (2). If $\Phi(t_{0})=0$ for some $t_{0}>0$, then $M^{\Phi}_{(\omega_{j})}(\mathbf{R}^{d})=M^{\infty}_{(\omega_{j})}(\mathbf{R}^{d})$, and the result follows from [60, Theorem 2.1]. Therefore suppose that $\Phi(t)>0$ when $t>0$. Then $\mathscr{S}(\mathbf{R}^{d})$ is dense in $M^{\Phi}_{(\omega_{j})}(\mathbf{R}^{d})$.
Suppose that $f_{1},f_{2}\in\mathscr{S}(\mathbf{R}^{d})$ satisfy
$\|f_{1}\|_{M^{\Phi}_{(\omega_{1})}}=\|f_{2}\|_{M^{\Psi}_{(\omega_{2})}}=1$ (2.18)
Then Lemma 2.7 gives
$|(\operatorname{Op}_{\varphi}(a)f_{1},f_{2})|$ $\leq\iiint_{\mathbf{R}^{4d+m}}(G*H_{a,\omega})(\Theta)|F_{1}(y,\eta)|\,|F_{2}(x,\xi)|\,dXd\xi d\eta,$
where
$H_{a,\omega}$ $=\sup_{X\in\mathbf{R}^{4d+m}}|V_{\phi}a(X,\,\cdot\,)\omega(X,\,\cdot\,)|$
and
$\Theta$ $=(\xi-\varphi^{\prime}_{x}(X),\eta-\varphi^{\prime}_{y}(X),-\varphi^{\prime}_{\zeta}(X)),\quad X=(x,y,\zeta).$
By letting $t_{1}=|F_{1}(y,\eta)|$ and $t_{2}=|F_{2}(x,\xi)|$ in $t_{1}t_{2}\leq\Phi(t_{1})+\Psi(t_{2})$, we obtain
$|(\operatorname{Op}_{\varphi}(a)f_{1},f_{2})|$ $\leq J_{1}+J_{2},$ (2.19)
where
$J_{1}$ $=\iiint_{\mathbf{R}^{4d+m}}(G*H_{a,\omega})(\Theta)\Phi(|F_{1}(y,\eta)|)\,dXd\xi d\eta$
and
$J_{2}$ $=\iiint_{\mathbf{R}^{4d+m}}(G*H_{a,\omega})(\Theta)\Psi(|F_{2}(x,\xi)|)\,dXd\xi d\eta$
We need to estimate $J_{1}$ and $J_{2}$. By taking $z=\varphi^{\prime}_{\zeta}(X)$, $\eta_{0}$, $y$, $\xi$ and $\eta$ as new variables of integrations, and using (0.2), it follows that
$J_{1}$ $\leq d\hskip-0.80002pt\bar{}\hskip 1.99997pt^{\,-1}\iiint_{\mathbf{R}^{4d+m}}(G*H_{a,\omega})(\xi-\kappa(y,z,\eta_{0}),\eta-\eta_{0},z)\Phi(|F_{1}(y,\eta)|)\,dy\,dz\,d\xi\,d\eta\,d\eta_{0}$
$=d\hskip-0.80002pt\bar{}\hskip 1.99997pt^{\,-1}\iiint_{\mathbf{R}^{4d+m}}(G*H_{a,\omega})(\xi,\eta-\eta_{0},z)\Phi(|F_{1}(y,\eta)|)\,dy\,dz\,d\xi\,d\eta\,d\eta_{0}$
$=d\hskip-0.80002pt\bar{}\hskip 1.99997pt^{\,-1}\|G*H_{a,\omega}\|_{L^{1}}\iint\Phi(|F_{1}(y,\eta)|)\,dyd\eta\leq Cd\hskip-0.80002pt\bar{}\hskip 1.99997pt^{\,-1}\|G*H_{a,\omega}\|_{L^{1}}$
for some continuous function $\kappa$ and constant $C>0$. In the last inequality we have used the fact that
$\iint\Phi(|F_{1}(y,\eta)|)\,dyd\eta\leq 1$
when $\|f_{1}\|_{M^{\Phi}_{(\omega_{1})}}=1$. It follows by Young’s inequality, (2.2) and Lemma 2.7 that
[…]
for some constant $C>0$. Hence
[…]
for some constant $C>0$.
If we instead take $x$, $z=\varphi^{\prime}_{\zeta}(X)$, $\xi$, $\eta$ and $\xi_{0}=\varphi^{\prime}_{x}(X)$ as new variables of integrations, it follows by similar arguments that
[…]
A combination of (2.19), (2.20) and (2.21) now gives
[…]
when (2.18) holds. The result now follows from this estimate, by homogeneity, duality and the fact that $\mathscr{S}(\mathbf{R}^{d})$ is dense in $M^{\Phi}_{(\omega_{1})}(\mathbf{R}^{d})$. ∎
Proof of Proposition 2.5.
Let $X=(x,y,\zeta)$,
$\vartheta(X,\xi,\eta,z)$ $=\omega(X,\xi,\eta,z)(1+|x|+|y|+|z|+|\xi|)^{3}$
and
$\vartheta_{2}(x,\xi)$ $=\omega_{2}(x,\xi)(1+|x|)(1+|\xi|).$
Since $|\varphi^{\prime\prime}(x,y,\zeta)|$ is a bounded function, it follows that
$|\varphi^{\prime}_{x}|=|\varphi^{\prime}_{x}(X)|\lesssim(1+|x|+|y|+|\zeta|).$
This gives
$\omega(X,\xi-\varphi^{\prime}_{x},\eta-\varphi^{\prime}_{y},-\varphi^{\prime}_{\zeta})(1+|x|)(1+|\xi|)$
$\leq\omega(X,\xi-\varphi^{\prime}_{x},\eta-\varphi^{\prime}_{y},-\varphi^{\prime}_{\zeta})(1+|x|)(1+|\xi-\varphi^{\prime}_{x}|)(1+|\varphi^{\prime}_{x}|)$
$\lesssim\omega(X,\xi-\varphi^{\prime}_{x},\eta-\varphi^{\prime}_{y},-\varphi^{\prime}_{\zeta})(1+|x|+|y|+|\zeta|+|\xi-\varphi^{\prime}_{x}|)^{3}$
$=\vartheta(X,\xi-\varphi^{\prime}_{x},\eta-\varphi^{\prime}_{y},-\varphi^{\prime}_{\zeta}).$
A combination of these estimates and (2.2) shows that (2.2) holds with $\vartheta$ and $\vartheta_{2}$ in place of $\vartheta$ and $\vartheta_{2}$.
Since $a\in\Sigma_{1}(\mathbf{R}^{2d+N})$, it follows that $a\in M^{\infty,1}_{(\vartheta)}(\mathbf{R}^{2d+N})$. Hence Theorem 2.3 shows that
$\operatorname{Op}_{\varphi}(a):M^{\Phi}_{(\omega_{1})}(\mathbf{R}^{d})\to M^{\Phi}_{(\vartheta_{2})}(\mathbf{R}^{d})$ (2.22)
is continuous.
Since the inclusion map
$\boldsymbol{\iota}:M^{\Phi}_{(\vartheta_{2})}(\mathbf{R}^{d})\to M^{\Phi}_{(\omega_{2})}(\mathbf{R}^{d}),$ (2.23)
is compact, in view of [61, Proposition 4.8], it follows that
$\operatorname{Op}_{\varphi}(a)=\boldsymbol{\iota}\circ\operatorname{Op}_{\varphi}(a):M^{\Phi}_{(\omega_{1})}(\mathbf{R}^{d})\to M^{\Phi}_{(\omega_{2})}(\mathbf{R}^{d})$
is compact, giving the result.
Proof of Theorem 2.4.
By Theorem 2.3 we have
$\|\operatorname{Op}_{\varphi}(a)\|_{M^{\Phi}_{(\omega_{1})}\to M^{\Phi}_{(\omega_{2})}}\lesssim\|a\|_{M^{\infty,1}_{(\omega)}}.$
Since $a\in M^{\sharp,1}_{(\omega)}(\mathbf{R}^{2d+N})$, it follows from Lemma 1.11 and Proposition 2.5, that for some $a_{j}\in\Sigma_{1}(\mathbf{R}^{2d+N})$, one has
$\|a-a_{j}\|_{M^{\infty,1}_{(\omega)}}\to 0,\quad\text{as}\quad j\to\infty.$
A combination of these estimates gives
$\|\operatorname{Op}_{\varphi}(a)-\operatorname{Op}_{\varphi}(a)\|_{M^{\Phi}_{(\omega_{1})}\to M^{\Phi}_{(\omega_{2})}}\lesssim\|a-a_{j}\|_{M^{\infty,1}_{(\omega)}}\to 0,\quad\text{as}\quad j\to\infty.$
Hence, $\operatorname{Op}_{\varphi}(a)$ can be approximated in norm by $\operatorname{Op}_{\varphi}(a_{j})$. Since $a_{j}\in\Sigma_{1}(\mathbf{R}^{2d+N})$, it follows that $\operatorname{Op}_{\varphi}(a_{j})$. A combination of these properties shows that $\operatorname{Op}_{\varphi}(a)$ is compact, and the result follows. ∎
Proof of Theorem 2.6.
The result follows from [60, Theorem 2.7] when $p_{\Phi}=q_{\Phi}$. Therefore we assume that $1<q_{\Phi}\leq p_{\Phi}<\infty$.
First we suppose that (i) holds. Let $\phi$, $\phi_{j}$ and $G$ be as in Lemma 2.7, and let
$U=|V_{\phi}a\cdot\omega|.$
By Lemma 2.7, Remark 2.9 and Minkowski’s inequality we obtain
$\|K_{a,\varphi}\|_{M^{\Phi}}$ $\lesssim\left\|\int_{\mathbf{R}^{m}}|\mathcal{K}_{a,\varphi}(X,\xi,\eta)|\,d\zeta\right\|_{L^{\Phi}}$
$\lesssim\left\|\int_{\mathbf{R}^{m}}(G*U(X,\,\cdot\,))(\Theta)\,d\zeta\right\|_{L^{\Phi}}$
$\leq\left\|\int_{\mathbf{R}^{m}}U_{2}(\,\cdot\,,\zeta)\,d\zeta\right\|_{L^{\Phi}},$ (2.24)
where
$U_{2}(X)$ $=U_{2}(x,y,\zeta)=\|U_{1}(X,\,\cdot\,-(\varphi^{\prime}_{x}(X),\varphi^{\prime}_{y}(X)),-\varphi^{\prime}_{\zeta}(X))\|_{L^{\Phi}}$
$=\|U_{1}(X,\,\cdot\,,-\varphi^{\prime}_{\zeta}(X))\|_{L^{\Phi}},$ (2.25)
with
$U_{1}(X,\xi,\eta,z)$ $=(G*U(X,\,\cdot\,))(\xi,\eta,z).$ (2.26)
By using Minkowski’s inequality again, we obtain
$\int_{\mathbf{R}^{m}}U_{2}(X)\,d\zeta$ $\leq\|G\|_{L^{1}}\int_{\mathbf{R}^{m}}\|U(X,\,\cdot\,,-\varphi^{\prime}_{\zeta}(X))\|_{L^{\Phi}}\,d\zeta$
$\leq\|G\|_{L^{1}}\int_{\mathbf{R}^{m}}\sup_{z\in\mathbf{R}^{m}}\left(\|U(X,\,\cdot\,,z)\|_{L^{\Phi}}\right)\,d\zeta.$
A combination of these estimates now gives (2.8), and thereby (1). The assertion (2) now follows from (1) and Proposition 2.1, and we have proved the result when (i) is fulfilled.
Next suppose that (ii) holds. Then for $U_{2}$ in (2.25) and (2.26) we have
$\int_{\mathbf{R}^{m}}U_{2}(x,y,\zeta)\,d\zeta$ $=\int_{\mathbf{R}^{m}}\|U_{1}(x,y,\zeta,\,\cdot\,,-\varphi^{\prime}_{\zeta}(X))\|_{L^{\Phi}}\,d\zeta$
$\leq\frac{1}{d\hskip-0.80002pt\bar{}\hskip 1.99997pt}\int_{\mathbf{R}^{m}}\|U_{1}(x,y,\zeta(x,y,z),\,\cdot\,,z)\|_{L^{\Phi}}\,dz$
$\leq\frac{1}{d\hskip-0.80002pt\bar{}\hskip 1.99997pt}\int_{\mathbf{R}^{m}}\sup_{\zeta\in\mathbf{R}^{m}}\left(\|U_{1}(x,y,\zeta,\,\cdot\,,z)\|_{L^{\Phi}}\right)\,dz.$
Here in, the first inequality we have taken $z=-\varphi^{\prime}_{\zeta}(X)$ as new variable of integration, and used that $|\det(\varphi^{\prime\prime}_{\zeta,\zeta})|\geq d\hskip-0.80002pt\bar{}\hskip 1.99997pt$.
A combination of these estimates Proposition 2.1 both (1) and (2) in the case when (ii) holds, giving the result. ∎
Remark 2.10.
Let $s>1$, $d=d_{2}+d_{1}$, $\Phi$ be a Young function such that $1<q_{\Phi}\leq p_{\Phi}<\infty$, $q_{\Phi}=p_{\Phi}=1$ or $q_{\Phi}=p_{\Phi}=\infty$, let $\phi\in\mathcal{D}_{0,s}(\mathbf{R}^{d+m})\setminus 0$, $\omega$, $\omega_{j}$, $j=0,1,2$, $v$ and $\varphi$ be as in (2.2) and (2.3). Also let $a\in\Sigma_{s}^{\prime}(\mathbf{R}^{d+m})$ be such that $\|a\|$ in (2.7) is finite, and that $|\det(\varphi^{\prime\prime}_{\zeta,\zeta})|\geq d\hskip-0.80002pt\bar{}\hskip 1.99997pt$. Then the following is true:
  1. (1)
    the kernel $K_{a,\varphi}$ of $\operatorname{Op}_{\varphi}(a)$ belongs to $M^{\Phi}_{(\omega_{0})}(\mathbf{R}^{d})$, and (2.8) holds for some constant $C$ which is independent of $a\in\mathscr{S}^{\prime}(\mathbf{R}^{d+m})$ and $\varphi\in C(\mathbf{R}^{d+m})$;
  2. (2)
    the definition of $\operatorname{Op}_{\varphi}(a)$ extends uniquely to a continuous operator from $M^{\Phi^{*}}_{(\omega_{1})}(\mathbf{R}^{d_{1}})$ to $M^{\Phi}_{(\omega_{2})}(\mathbf{R}^{d_{2}})$, and (2.9) holds.
3. Schatten-von Neumann properties for Fourier integral operators
In this section we discuss Orlicz Schatten-von Neumann properties for Fourier integral operators with amplitudes, essentially only depending on two variables, belonging to Orlicz modulation spaces. In the first part we consider Fourier integral operators of the form
$\operatorname{Op}_{\varphi}(a)f(x)$ $=(2\pi)^{-d}\iint_{\mathbf{R}^{2d}}a(x,\zeta)f(y)e^{i\varphi(x,y,\zeta)}\,dyd\zeta.$ (3.1)
Thereafter we perform some extensions to Fourier integral operators of the form
$\operatorname{Op}_{A,\varphi}(a)f(x)$ $=(2\pi)^{-d}\iint_{\mathbf{R}^{2d}}a(x-A(x-y),\zeta)f(y)e^{i\varphi(x,y,\zeta)}\,dyd\zeta.$ (3.1)
Here $A$ can be any real $d\times d$ matrix. (See also (0.4) and (0.5) from the introduction.) In the passage from the operators in (3.1) to (3.1) it is sometimes convenient to also let the phase function $\varphi$ in (3.1) depend on $A$ as
$\operatorname{Op}_{A,\varphi}(a)f(x)$ $=(2\pi)^{-d}\iint_{\mathbf{R}^{2d}}a(x-A(x-y),\zeta)f(y)e^{i\varphi_{A}(x,y,\zeta)}\,dyd\zeta,$ (3.1)′′
with
$\varphi_{A}(x,y,\zeta)$ $=\varphi(x-A(x-y),y-A(x-y),\zeta).$ (3.2)
We observe that the map which takes amplitudes to the kernels of these operators are formally given by
$a$ $\mapsto K_{a,\varphi}(x,y)\equiv(2\pi)^{-d}\int_{\mathbf{R}^{d}}a(x,\zeta)e^{i\varphi(x,y,\zeta)}\,d\zeta,$ (3.3)
$a$ $\mapsto K_{a,\varphi,A}(x,y)\equiv(2\pi)^{-d}\int_{\mathbf{R}^{d}}a(x-A(x-y),\zeta)e^{i\varphi(x,y,\zeta)}\,d\zeta,$ (3.3)
and
$a$ $\mapsto K_{a,\varphi,A}(x,y)\equiv(2\pi)^{-d}\int_{\mathbf{R}^{d}}a(x-A(x-y),\zeta)e^{i\varphi_{A}(x,y,\zeta)}\,d\zeta,$ (3.3)′′
provided we may interpret the integrals in some sense.
Our first result on this is the following. Here we assume that the involved weight and phase functions satisfies
$\omega_{0}(x,y,\xi+\varphi^{\prime}_{x}(X),\varphi^{\prime}_{y}(X))$ $\leq C\omega(x,\zeta,\xi,-\varphi^{\prime}_{\zeta}(X)),$ (3.4)
$\frac{\omega_{2}(x,\xi)}{\omega_{1}(y,\eta)}$ $\leq C\omega_{0}(x,y,\xi,-\eta)$ (3.5)
$\omega_{0}(x,y,\xi,\eta_{1}+\eta_{2})$ $\leq C\omega_{0}(x,y,\xi,\eta_{1})v_{1}(\eta_{2})$ (3.6)
$\omega(x,\zeta,\xi_{1}+\xi_{2},z_{1}+z_{2})$ $\leq\omega(x,\zeta,\xi_{1},z_{1})v_{2}(\xi_{2},z_{2}),$ (3.7)
$v(X,\xi,\eta,z)$ $=v_{1}(\eta)v_{2}(\xi,z),$ (3.8)
$x,y,z,z_{j},\xi,\xi_{j},\eta,\eta_{j},\zeta$ $\in\mathbf{R}^{d},\quad X=(x,y,\zeta),\quad j=1,2.$
Proposition 3.1.
Suppose $\Phi$ is a Young function which satisfies a local $\Delta_{2}$ condition, $s>1$, $d\hskip-0.80002pt\bar{}\hskip 1.99997pt>0$, $\omega,\omega_{0}\in\mathscr{P}_{s}(\mathbf{R}^{4d})$ and $v\in\mathscr{P}_{s}(\mathbf{R}^{6d})$ satisfy (3.4), (3.6)–(3.8), and let $\varphi\in C^{2}(\mathbf{R}^{3d})$ be such that $\varphi^{\prime\prime}\in M^{\infty,1}_{(v)}(\mathbf{R}^{3d})$ and
$|\det(\varphi^{\prime\prime}_{y,\zeta})|\geq d\hskip-0.80002pt\bar{}\hskip 1.99997pt.$ (3.9)
Then the map in (3.3) from $\Sigma_{1}(\mathbf{R}^{2d})$ to $\Sigma_{1}^{\prime}(\mathbf{R}^{2d})$ extends uniquely to a continuous map from $M^{\Phi}_{(\omega)}(\mathbf{R}^{2d})$ to $M^{\Phi}_{(\omega_{0})}(\mathbf{R}^{2d})$, and
$\|K_{a,\varphi}\|_{M^{\Phi}_{(\omega_{0})}}\leq Cd\hskip-0.80002pt\bar{}\hskip 1.99997pt^{-1}\exp(\|\varphi\|_{M^{\infty,1}_{(v)}})\|a\|_{M^{\Phi}_{(\omega)}},\qquad a\in M^{\Phi}_{(\omega)}(\mathbf{R}^{2d}),$ (3.10)
for some constant $C$ which is independent of $a$, $\varphi$ and $\Phi$.
For the proof we observe that $M^{\infty,1}_{(v)}(\mathbf{R}^{3d})\subseteq C(\mathbf{R}^{3d})\cap L^{\infty}(\mathbf{R}^{3d})$. Hence the conditions on $\varphi$ in Proposition 3.1 imply that
$d\hskip-0.80002pt\bar{}\hskip 1.99997pt\leq|\det(\varphi^{\prime\prime}_{y,\zeta}(x,y,\zeta))|\leq C,$ (3.11)
for some constant $C>0$ which is independent of $x,y,\zeta\in\mathbf{R}^{d}$.
Proof of Proposition 3.1.
Let $\phi$ and $\phi_{j}$ be the same as in (2.13), where we additionally assume that $\phi$ is given by
$\phi(x,y,\zeta)=\psi_{1}(x,\zeta)\psi_{2}(y),\qquad x,y,\zeta\in\mathbf{R}^{d}.$
Also let $\psi$ be as in Remark 2.9, and let
$H=|V_{\psi_{1}}a\cdot\omega|.$
Then $\|a\|_{M^{\Phi}_{(\omega)}}=\|H\|_{L^{\Phi}}$.
If $\Xi_{1}=(\xi_{1},\eta_{1},z_{1})\in\mathbf{R}^{3d}$, then Lemma 2.7 and Remark 2.9 give
$|V_{\psi}K_{a,\varphi}(x,y,\xi,\eta)\cdot\omega_{0}(x,y,\xi,\eta)|\asymp\left|\int_{\mathbf{R}^{d}}\mathcal{K}_{a,\varphi}(X,\xi,\eta)\,d\zeta\right|$ (3.12)
$\leq\iint\limits_{\mathbf{R}^{4d}}H(x,\zeta,\xi-\varphi^{\prime}_{x}-\xi_{1},-\varphi^{\prime}_{\zeta}-z_{1})|\widehat{\psi}_{2}(\eta-\varphi^{\prime}_{y}-\eta_{1})|G(\Xi_{1})\,d\Xi_{1}d\zeta.$
Here observe that
$(\varphi^{\prime}_{x},\varphi^{\prime}_{y},\varphi^{\prime}_{\zeta})=(\varphi^{\prime}_{x}(X),\varphi^{\prime}_{y}(X),\varphi^{\prime}_{\zeta}(X))$
depends on $X=(x,y,\zeta)\in\mathbf{R}^{3d}$. In order to deduce the result we shall consider two cases. In the first case we assume that $\Phi$ is positive, and in the second case we assume that $\Phi$ fails to be positive.
Therefore, suppose that $\Phi$ additionally is positive. Then by replacing $\Phi$ with another Young function which agree with $\Phi$ near origin, we may assume that $\Phi$ satisfies a global $\Delta_{2}$-condition.
By taking $(\Xi_{1},\varphi^{\prime}_{y})$ as new variables of integrations, and using (3.11) and (3.12), we obtain
$|V_{\psi}K_{a,\varphi}(x,y,\xi,\eta)\cdot\omega_{0}(x,y,\xi,\eta)|$
$\leq\frac{1}{d\hskip-0.80002pt\bar{}\hskip 1.99997pt}\iint\limits_{\mathbf{R}^{4d}}H(x,\zeta,\xi-\varphi^{\prime}_{x}-\xi_{1},-\varphi^{\prime}_{\zeta}-z_{1})|\widehat{\psi}_{2}(\eta-\zeta_{1}-\eta_{1})|G(\Xi_{1})\,d\Xi_{1}d\zeta_{1}.$
Let $C_{0}=\|G\|_{L^{1}}\|\widehat{\psi}_{2}\|_{L^{1}}$. Then Jensen’s inequality gives
$\Phi(|V_{\psi}K_{a,\varphi}(x,y,\xi,\eta)\cdot\omega_{0}(x,y,\xi,\eta)|)$
$\leq\Phi\left(\frac{1}{d\hskip-0.80002pt\bar{}\hskip 1.99997pt}\iint\limits_{\mathbf{R}^{4d}}H(x,\zeta,\xi-\varphi^{\prime}_{x}-\xi_{1},-\varphi^{\prime}_{\zeta}-z_{1})|\widehat{\psi}_{2}(\eta-\zeta_{1}-\eta_{1})|G(\Xi_{1})\,d\Xi_{1}d\zeta_{1}\right)$
$\leq\frac{1}{C_{0}}\iint\limits_{\mathbf{R}^{4d}}\Phi\left(\frac{C_{0}}{d\hskip-0.80002pt\bar{}\hskip 1.99997pt}H(x,\zeta,\xi-\varphi^{\prime}_{x}-\xi_{1},-\varphi^{\prime}_{\zeta}-z_{1})\right)|\widehat{\psi}_{2}(\eta-\zeta_{1}-\eta_{1})|G(\Xi_{1})\,d\Xi_{1}d\zeta_{1}$
$\leq\frac{C}{C_{0}}\iint\limits_{\mathbf{R}^{4d}}\Phi\left(\frac{C_{0}}{d\hskip-0.80002pt\bar{}\hskip 1.99997pt}H(x,\zeta,\xi-\varphi^{\prime}_{x}-\xi_{1},-\varphi^{\prime}_{\zeta}-z_{1})\right)|\widehat{\psi}_{2}(\eta-\varphi^{\prime}_{y}-\eta_{1})|G(\Xi_{1})\,d\Xi_{1}d\zeta.$
In the last step we have taken back the original variables of integration, and used the second inequality in (3.11). Since $\Phi$ satisfies a $\Delta_{2}$ condition we get
$\Phi(|V_{\psi}K_{a,\varphi}(x,y,\xi,\eta)\cdot\omega_{0}(x,y,\xi,\eta)|)$
$\leq C\iint\limits_{\mathbf{R}^{4d}}\Phi\left(H(x,\zeta,\xi-\varphi^{\prime}_{x}-\xi_{1},-\varphi^{\prime}_{\zeta}-z_{1})\right)|\widehat{\psi}_{2}(\eta-\varphi^{\prime}_{y}-\eta_{1})|G(\Xi_{1})\,d\Xi_{1}d\zeta,$
for some constant $C>0$.
By integration we get
$\iiiint\limits_{\mathbf{R}^{4d}}\Phi(|V_{\psi}K_{a,\varphi}(x,y,\xi,\eta)\cdot\omega_{0}(x,y,\xi,\eta)|)\,dxdyd\xi d\eta$
$\leq C\iiiint\limits_{\mathbf{R}^{8d}}\Phi\left(H(x,\zeta,\xi-\varphi^{\prime}_{x}-\xi_{1},-\varphi^{\prime}_{\zeta}-z_{1})\right)|\widehat{\psi}_{2}(\eta-\varphi^{\prime}_{y}-\eta_{1})|G(\Xi_{1})\,dXd\xi d\eta d\Xi_{1}.$
Here recall that $X=(x,y,\zeta)$, giving that the integration variables in the last integral are given by
$(x,y,\zeta,\xi,\eta,\xi_{1},\eta_{1},z_{1})\in\mathbf{R}^{8d}.$
By taking
$(x,\varphi^{\prime}_{\zeta},\zeta,\xi,\eta,\xi_{1},\eta_{1},z_{1})$
as new variables of integrations, and using the first inequality in (3.11), we obtain
$\iiiint\limits_{\mathbf{R}^{4d}}\Phi(|V_{\psi}K_{a,\varphi}(x,y,\xi,\eta)\cdot\omega_{0}(x,y,\xi,\eta)|)\,dxdyd\xi d\eta$
$\leq\frac{C}{d\hskip-0.80002pt\bar{}\hskip 1.99997pt}\iiiint\limits_{\mathbf{R}^{8d}}\Phi\left(H(x,\zeta,\xi-\varphi^{\prime}_{x}-\xi_{1},-y-z_{1})\right)|\widehat{\psi}_{2}(\eta-\varphi^{\prime}_{y}-\eta_{1})|G(\Xi_{1})\,dXd\xi d\eta d\Xi_{1}$
$=\frac{C}{d\hskip-0.80002pt\bar{}\hskip 1.99997pt}\iiiint\limits_{\mathbf{R}^{8d}}\Phi\left(H(x,\zeta,\xi,y)\right)|\widehat{\psi}_{2}(\eta)|G(\Xi_{1})\,dXd\xi d\eta d\Xi_{1}$
$=\frac{C}{d\hskip-0.80002pt\bar{}\hskip 1.99997pt}\|G\|_{L^{1}}\|\psi_{2}\|_{L^{1}}\|\Phi(H)\|_{L^{1}}.$
Here in the first equality we have performed a straight-forward substitution of integration variables.
The estimate (3.10) now follows from the last estimates, homogeneity and the fact that $\Phi$ satisfies a $\Delta_{2}$ condition. This gives the result when $\Phi$ is positive.
Next suppose that $\Phi$ is not positive. Then $\Phi$ equals to zero near origin, which implies that $M^{\Phi}_{(\omega)}=M^{\infty}_{(\omega)}$ and $M^{\Phi}_{(\omega_{0})}=M^{\infty}_{(\omega_{0})}$. By (3.10) we obtain
$\|K_{a,\varphi}\|_{M^{\infty}_{(\omega_{0})}}$ $=\sup_{x,y,\xi,\eta}\big(|V_{\psi}K_{a,\varphi}(x,y,\xi,\eta)\cdot\omega_{0}(x,y,\xi,\eta)|\big)$
$\lesssim\|H\|_{L^{\infty}}\sup_{x,y,\xi,\eta}\left(\iint\limits_{\mathbf{R}^{4d}}|\widehat{\psi}_{2}(\eta-\varphi^{\prime}_{y}-\eta_{1})|G(\Xi_{1})\,d\Xi_{1}d\zeta\right)$
$\leq\|H\|_{L^{\infty}}\sup_{x,y,\xi,\eta}\left(\frac{1}{d\hskip-0.80002pt\bar{}\hskip 1.99997pt}\iint\limits_{\mathbf{R}^{4d}}|\widehat{\psi}_{2}(\eta-\zeta_{1}-\eta_{1})|G(\Xi_{1})\,d\Xi_{1}d\zeta_{1}\right)$
$\leq\|H\|_{L^{\infty}}\asymp\|a\|_{M^{\infty}_{(\omega)}}.$
In the second inequality we have again taking $(\Xi_{1},\varphi^{\prime}_{y})$ as new variables of integration. The result now follows in this case from the latter estimates. ∎
We shall combine the previous proposition with the following one.
Proposition 3.2.
Let $\Phi$ be a quasi-Young function which satisfies
$q_{\Phi}\leq p_{\Phi}<2\quad\text{or}\quad q_{\Phi}=p_{\Phi}=2.$
Also let $\omega_{0}\in\mathscr{P}_{E}(\mathbf{R}^{2d_{2}+2d_{1}})$ and $\omega_{j}\in\mathscr{P}_{E}(\mathbf{R}^{2d_{j}})$, $j=1,2$, be such that (3.5) holds. If $K\in M^{\Phi}_{(\omega)}(\mathbf{R}^{d_{2}+d_{1}})$, then $T_{K}\in\mathscr{I}_{\Phi}(\omega_{1},\omega_{2})$, and
$\|T_{K}\|_{\mathscr{I}_{\Phi}(\omega_{1},\omega_{2})}\leq C\|K\|_{M^{\Phi}_{(\omega_{0})}},\qquad K\in M^{\Phi}_{(\omega)}(\mathbf{R}^{d_{2}+d_{1}}),$
for some constant $C>0$ which is independent of $K$ and $\Phi$.
Proof.
Let $d=d_{2}+d_{1}$, $\phi,\psi\in\Sigma_{1}(\mathbf{R}^{d})$ and $\varepsilon>0$ be chosen such that
$\{\phi_{j,\iota}\}_{j,\iota\in\varepsilon\mathbf{Z}^{d}}$ and $\{\psi_{j,\iota}\}_{j,\iota\in\varepsilon\mathbf{Z}^{d}},$
$\phi_{j,\iota}=e^{i\langle\,\cdot\,,\iota\rangle}\phi(\,\cdot\,-j),$ $\psi_{j,\iota}=e^{i\langle\,\cdot\,,\iota\rangle}\psi(\,\cdot\,-j),$
are dual Gabor frames. The existence of such frames follows from e. g. [30].
For any sequence $c=\{c(j,\iota)\}_{j,\iota\in\varepsilon\mathbf{Z}^{2d}}$, let $T_{c}$ be the kernel operator
$T_{c}=T_{K}=\sum_{j,\iota}c(j,\iota)T_{\phi_{j,\iota}}\quad\text{when}\quad K=K_{c}=\sum_{j,\iota}c(j,\iota)\phi_{j,\iota}$ (3.13)
(i. e. $K_{c}$ is the synthesis of $c$). Also let $\operatorname{ON}(\omega_{j})$ denote the set of all orthonormal sequences in $M^{2}_{(\omega_{j})}(\mathbf{R}^{d_{j}})$, $\{f_{j,k}\}_{k=1}^{\infty}\in\operatorname{ON}(\omega_{j})$, $j=1,2$, and let $S$ be the operator from $\ell_{0}(\varepsilon\mathbf{Z}^{2d})$ to $\ell_{0}^{\prime}(\mathbf{Z}_{+})$, given by
$S(\{c(j,\iota)\}_{j,\iota\in\varepsilon\mathbf{Z}^{2d}})=\{(T_{c}f_{1,k},f_{2,k})_{M^{2}_{(\omega_{2})}}\}_{k=1}^{\infty}.$
We have
$\|\{c(j,\iota)\}_{j,\iota\in\varepsilon\mathbf{Z}^{2d}}\|_{\ell^{\Phi}_{(\omega)}}$ $\asymp\|K_{c}\|_{M^{\Phi}_{(\omega)}}$ (3.14)
when $c\in\ell_{0}(\varepsilon\mathbf{Z}^{2d})$, and
$\|\{(T_{c}f_{1,k},f_{2,k})_{M^{2}_{(\omega_{2})}}\}_{k=1}^{\infty}\|_{\ell^{\Psi}}$ $\leq\|T_{c}\|_{\mathscr{I}_{\Psi}(\omega_{1},\omega_{2})}.$ (3.15)
We have
$\|T_{K}\|_{\mathscr{I}_{p}(\omega_{1},\omega_{2})}\leq C\|K\|_{M^{p}_{(\omega)}},$
for some constant $C>0$ which is independent of $K\in M^{p}_{(\omega)}(\mathbf{R}^{d})$ and $0<p\leq 2$ (see e. g. [59]). Hence it follows by combining these relations that
$\|S(c)\|_{\ell^{p}}\leq C\|c\|_{\ell^{p}_{(\omega)}},$ (3.16)
for some constant $C>0$ which is independent of $c\in\ell_{0}(\varepsilon\mathbf{Z}^{2d})$, $\{f_{j,k}\}_{k=1}^{\infty}\in\operatorname{ON}(\omega_{j})$, $j=1,2$, and $0<p\leq 2$. Since $\ell_{0}$ is dense in $\ell^{p}$ when $0<p\leq 2$, it follows that $S$ extends uniquely to a continuous map from $\ell^{p}_{(\omega)}(\varepsilon\mathbf{Z}^{2d})$ to $\ell^{p}(\mathbf{Z}_{+})$, and that (3.16) holds for any $c\in\ell^{p}_{(\omega)}(\varepsilon\mathbf{Z}^{2d})$.
By Marcinkiewicz interpolation theorem for Orlicz spaces, given in [37], it follows that $S$ restricts to a continuous map from $\ell^{\Phi}_{(\omega)}(\varepsilon\mathbf{Z}^{2d})$ to $\ell^{\Phi}(\mathbf{Z}_{+})$, and that
$\|S(c)\|_{\ell^{\Phi}}\leq C\|c\|_{\ell^{\Phi}_{(\omega)}},\qquad c\in\ell^{\Phi}_{(\omega)}(\varepsilon\mathbf{Z}^{2d}).$ (3.16)
A combination of (3.14), (3.15) and (3.16) gives
$\|\{(T_{c}f_{1,k},f_{2,k})_{M^{2}_{(\omega_{2})}}\}_{k=1}^{\infty}\|_{\ell^{\Psi}}\leq C\|K_{c}\|_{M^{p}_{(\omega)}},\qquad c\in\ell^{\Phi}_{(\omega)}(\varepsilon\mathbf{Z}^{2d}).$
Since any $K\in M^{\Phi}_{(\omega)}(\mathbf{R}^{d})$ is given by $K_{c}$, for some $c\in\ell^{\Phi}_{(\omega)}(\varepsilon\mathbf{Z}^{2d})$, the last estimate gives
$\|\{(T_{K}f_{1,k},f_{2,k})_{M^{2}_{(\omega_{2})}}\}_{k=1}^{\infty}\|_{\ell^{\Psi}}\leq C\|K_{c}\|_{M^{p}_{(\omega)}},\qquad K\in M^{\Phi}_{(\omega)}(\mathbf{R}^{d}),$
where we recall that the constant $C$ is independent of the choices of $\{f_{j,k}\}_{k=1}^{\infty}\in\operatorname{ON}(\omega_{j})$, $j=1,2$. By taking the supremum over all such orthonormal sequences we obtain
$\|T_{K}\|_{\mathscr{I}_{\Phi}(\omega_{1},\omega_{2})}\leq C\|K_{c}\|_{M^{p}_{(\omega)}},\qquad K\in M^{\Phi}_{(\omega)}(\mathbf{R}^{d}),$
which gives the result. ∎
A straight-forward combination of Propositions 3.1 and 3.2 gives the following. The details are left for the reader.
Theorem 3.3.
Suppose $\Phi$ is a Young function which satisfies a local $\Delta_{2}$ condition, either $1<q_{\Phi}\leq p_{\Phi}<2$ or $1\leq q_{\Phi}=p_{\Phi}\leq 2$, $s>1$, $d\hskip-0.80002pt\bar{}\hskip 1.99997pt>0$, $\omega,\omega_{0}\in\mathscr{P}_{s}(\mathbf{R}^{4d})$ and $v\in\mathscr{P}_{s}(\mathbf{R}^{6d})$ satisfy (3.4), (3.5)–(3.8), $\varphi\in C^{2}(\mathbf{R}^{3d})$ satisfies $\varphi^{\prime\prime}\in M^{\infty,1}_{(v)}(\mathbf{R}^{3d})$ and (3.9), and let $a\in M^{\Phi}_{(\omega)}(\mathbf{R}^{2d})$. Then $\operatorname{Op}_{\varphi}(a)\in\mathscr{I}_{\Phi}(\omega_{1},\omega_{2})$, and
$\|\operatorname{Op}_{\varphi}(a)\|_{\mathscr{I}_{\Phi}(\omega_{1},\omega_{2})}$ $\leq Cd\hskip-0.80002pt\bar{}\hskip 1.99997pt^{-1}\exp(\|\varphi\|_{M^{\infty,1}_{(v)}})\|a\|_{M^{\Phi}_{(\omega)}},$ (3.17)
$a$ $\in M^{\Phi}_{(\omega)}(\mathbf{R}^{2d}),$
for some constant $C$ which is independent of $a$, $\varphi$ and $\Phi$.
Next we extend the previous results and investigations to include Fourier integral operators of the form (3.1). The following proposition is important for this transition. Here for any weight function $\omega_{0}$ on $\mathbf{R}^{2d}$ and real $d\times d$ matrix $A$, we let
$\omega_{A}(x,y,\xi,\eta)$
$\equiv\omega_{0}(x-A(x-y),y-A(x-y),\xi+A^{*}(\xi+\eta),\eta-A^{*}(\xi+\eta)).$ (3.18)
It follows that
$\omega_{0}(x,y,\xi,\eta)$
$=\omega_{A}(x+A(x-y),y+A(x-y),\xi-A^{*}(\xi+\eta),\eta+A^{*}(\xi+\eta)).$
Here also recall that for a linear bijective map $T$ on $\mathbf{R}^{d}$, the pullback $T^{*}f$ of $T$ on a distribution $f$ on $\mathbf{R}^{d}$ is defined by the formula
$\langle T^{*}f,\phi\rangle=|\det(T)|^{-1}\langle f,\phi(T^{-1}\,\cdot\,)\rangle,$
for any test function $\phi$ in corresponding test function space.
Proposition 3.4.
Let $A$ be a real $d\times d$ matrix, $\Phi$ be a Young function, and let $\omega_{0},\omega_{A}\in\mathscr{P}_{E}(\mathbf{R}^{2d})$ be such that (3.18) holds. Also let $T_{A}:\mathbf{R}^{2d}\to\mathbf{R}^{2d}$ be given by
$T_{A}(x,y)\equiv(x-A(x-y),y-A(x-y)).$
Then $T_{A}^{*}$ on $\Sigma_{1}(\mathbf{R}^{2d})$ extends uniquely to a homoeomorphism from $M^{\Phi}_{(\omega_{0})}(\mathbf{R}^{2d})$ to $M^{\Phi}_{(\omega_{A})}(\mathbf{R}^{2d})$.
Proof.
Let $K\in\Sigma_{1}^{\prime}(\mathbf{R}^{2d})$, $\phi\in\Sigma_{1}(\mathbf{R}^{2d})\setminus 0$, $K_{A}=T_{A}^{*}K$ and $\phi_{A}=T_{A}^{*}\phi$. Then $\phi_{A}\in\Sigma_{1}(\mathbf{R}^{2d})\setminus 0$. By straight-forward computations it follows that
$V_{\phi_{A}}K_{A}(x,y,\xi,\eta)=V_{\phi}K(x-A(x-y),x-A(x-y),\xi+A^{*}(\xi+\eta),\eta-A^{*}(\xi+\eta)).$
By multiplying the equality by $\omega_{A}$ in (3.18) and applying the $L^{\Phi}$ norm we get
$\|V_{\phi_{A}}K_{A}\cdot\omega_{A}\|_{L^{\Phi}}=\|V_{\phi_{A}}K\cdot\omega_{0}\|_{L^{\Phi}}.$ (3.19)
In the last equality we have used the fact that
$\det\left(\begin{matrix}I-A&A\\ -A&I+A\end{matrix}\right)=\det\left(\begin{matrix}I+A^{*}&A^{*}\\ -A^{*}&I-A^{*}\end{matrix}\right)=1.$
The assertion is now a straight-forward consequence of (3.19). ∎
Remark 3.5.
In [61], a broad family of modulation spaces are presented, where each modulation space $M(\omega,\mathscr{B})$ is parameterized by a weight $\omega\in\mathscr{P}_{E}(\mathbf{R}^{2d})$ and normal invariant quasi-Banach function space $\mathscr{B}$ on $\mathbf{R}^{2d}$. If $\omega_{0}\in\mathscr{P}_{E}(\mathbf{R}^{4d})$ and $\mathscr{B}$ hosts functions defined on $\mathbf{R}^{4d}$, then the previous proof shows that $T^{*}A$ in Proposition 3.4 is a homeomorphism from $M(\omega_{0},\mathscr{B})$ to $M(\omega_{A},\mathscr{B})$. This extends Proposition 3.4 to more general modulation spaces.
Next we extend Proposition 3.1 to Fourier integral operators of the form (3.1). Here the conditions of involved weight functions need to be modified into
$\omega_{0}(x+A(x-y),y+A(x-y),(I-A^{*})\xi+\varphi^{\prime}_{x}(X),A^{*}\xi+\varphi^{\prime}_{y}(X))\\[4.30554pt] \leq C\omega(x,\zeta,\xi,-\varphi^{\prime}_{\zeta}(X)),\quad X=(x,y,\zeta).$ (3.19)
Proposition 3.1.
Suppose $\Phi$ is a Young function which satisfies a local $\Delta_{2}$ condition, $s>1$, $d\hskip-0.80002pt\bar{}\hskip 1.99997pt>0$, $\omega,\omega_{0}\in\mathscr{P}_{s}(\mathbf{R}^{4d})$ and $v\in\mathscr{P}_{s}(\mathbf{R}^{6d})$ satisfy 3.19, (3.6)–(3.8), and let $\varphi\in C^{2}(\mathbf{R}^{3d})$ be such that $\varphi^{\prime\prime}\in M^{\infty,1}_{(v)}(\mathbf{R}^{3d})$ and
$\left|\det\left(\varphi^{\prime\prime}_{y,\zeta}-A^{*}(\varphi^{\prime\prime}_{x,\zeta}+\varphi^{\prime\prime}_{y,\zeta})\right)\right|\geq d\hskip-0.80002pt\bar{}\hskip 1.99997pt.$ (3.9)
Then the map in (3.3) from $\Sigma_{1}(\mathbf{R}^{2d})$ to $\Sigma_{1}^{\prime}(\mathbf{R}^{2d})$ extends uniquely to a continuous map from $M^{\Phi}_{(\omega)}(\mathbf{R}^{2d})$ to $M^{\Phi}_{(\omega_{0})}(\mathbf{R}^{2d})$, and
$\|K_{a,\varphi,A}\|_{M^{\Phi}_{(\omega_{0})}}\leq Cd\hskip-0.80002pt\bar{}\hskip 1.99997pt^{-1}\exp(\|\varphi\|_{M^{\infty,1}_{(v)}})\|a\|_{M^{\Phi}_{(\omega)}},\qquad a\in M^{\Phi}_{(\omega)}(\mathbf{R}^{2d}),$ (3.10)
for some constant $C$ which is independent of $a$, $\varphi$ and $\Phi$.
Proof.
We shall reduce ourselves to the case when $A=0$, and then Proposition 3.1 will give the result.
Let
$\psi(x,y,\zeta)$ $\equiv\varphi(x+A(x-y),y+A(x-y),\zeta)$
and
$\vartheta(x,y,\xi,\eta)$ $\equiv\omega_{0}(x+A(x-y),y+A(x-y),\xi-A^{*}(\xi+\eta),\eta+A^{*}(\xi+\eta)).$
Then
$\varphi(x,y,\zeta)$ $=\psi(x-A(x-y),y-A(x-y),\zeta)$
and
$\omega_{0}(x,y,\xi,\eta)$ $=\vartheta(x-A(x-y),y-A(x-y),\xi+A^{*}(\xi+\eta),\eta-A^{*}(\xi+\eta)).$
Furthermore, the change rule gives
$\psi^{\prime}_{x}$ $=\varphi^{\prime}_{x}+A^{*}(\varphi^{\prime}_{x}+\varphi^{\prime}_{y}),\quad\psi^{\prime}_{y}=\varphi^{\prime}_{y}+A^{*}(\varphi^{\prime}_{x}+\varphi^{\prime}_{y}),\quad\psi^{\prime}_{\zeta}=\varphi^{\prime}_{\zeta}$ (3.20)
$\varphi^{\prime}_{x}$ $=\psi^{\prime}_{x}-A^{*}(\varphi^{\prime}_{x}+\varphi^{\prime}_{y}),\quad\varphi^{\prime}_{y}=\psi^{\prime}_{y}-A^{*}(\varphi^{\prime}_{x}+\varphi^{\prime}_{y}),$ (3.21)
and
$\psi_{y,\zeta}^{\prime\prime}$ $=\varphi_{y,\zeta}^{\prime\prime}-A^{*}(\varphi_{x,\zeta}^{\prime\prime}+\varphi_{y,\zeta}^{\prime\prime}).$ (3.22)
By (3.9) and (3.22) we get $|\det(\psi^{\prime\prime}_{y,\zeta})|\geq d\hskip-0.80002pt\bar{}\hskip 1.99997pt$. In particular, (3.9) holds with $\psi$ in place of $\varphi$.
Let $z_{1}=x+A(x-y)$ and $z_{2}=y+A(x-y)$. A combination of 3.19 and (3.20) also gives
$\vartheta(x,y,\xi+\psi^{\prime}_{x},\psi^{\prime}_{y})$
$=\omega_{0}(z_{1},z_{2},\xi+\psi^{\prime}_{x}-A^{*}(\xi+\psi^{\prime}_{x}+\psi^{\prime}_{y}),\psi^{\prime}_{y}+A^{*}(\xi+\psi^{\prime}_{x}+\psi^{\prime}_{y}))$
$=\omega_{0}(z_{1},z_{2},(I-A^{*})\xi+\varphi^{\prime}_{x},A^{*}\xi+\varphi^{\prime}_{y})$
$=\omega_{0}(x+A(x-y),y+A(x-y),(I-A^{*})\xi+\varphi^{\prime}_{x},A^{*}\xi+\varphi^{\prime}_{y}).$
$\leq C\omega(x,\zeta,\xi,-\varphi^{\prime}_{\zeta})=C\omega(x,\zeta,\xi,-\psi^{\prime}_{\zeta}),$
which shows that (3.4) holds with $\psi$ in place of $\varphi$.
Consequently, all assumptions in Proposition 3.1 are fulfilled with $\psi$ and $\vartheta$ in place of $\varphi$ and $\omega_{0}$, respectively. Hence Proposition 3.1 and in particular (3.10) give $K_{a,\psi}\in M^{\Phi}_{(\vartheta)}$, and that
$\|K_{a,\psi}\|_{M^{\Phi}_{(\vartheta)}}\leq Cd\hskip-0.80002pt\bar{}\hskip 1.99997pt^{-1}\exp(\|\varphi\|_{M^{\infty,1}_{(v)}})\|a\|_{M^{\Phi}_{(\omega)}},\qquad a\in M^{\Phi}_{(\omega)}(\mathbf{R}^{2d}).$ (3.10)′′
Since
$K_{a,\varphi,A}(x,y)=K_{a,\psi}(x-A(x-y),y-A(x-y)),$
it follows from Proposition 3.4 that
$\|K_{a,\varphi,A}\|_{M^{\Phi}_{(\omega_{0})}}\asymp\|K_{a,\psi}\|_{M^{\Phi}_{(\vartheta)}}.$
The estimate (3.10) now follows by combining the last relation with (3.10)′′, and the result follows. ∎
In similar ways as for Theorem 3.3 and its proof, we get the following extension by combining Proposition 3.2 with Proposition 3.1. The details are left for the reader.
Theorem 3.3.
Suppose AA is a real $d\times d$ matrix, $\Phi$ is a Young function which satisfies a local $\Delta_{2}$ condition, either $1<q_{\Phi}\leq p_{\Phi}<2$ or $1\leq q_{\Phi}=p_{\Phi}\leq 2$, $s>1$, $d\hskip-0.80002pt\bar{}\hskip 1.99997pt>0$, $\omega,\omega_{0}\in\mathscr{P}_{s}(\mathbf{R}^{4d})$ and $v\in\mathscr{P}_{s}(\mathbf{R}^{6d})$ satisfy 3.19, (3.5)–(3.8), $\varphi\in C^{2}(\mathbf{R}^{3d})$ satisfies $\varphi^{\prime\prime}\in M^{\infty,1}_{(v)}(\mathbf{R}^{3d})$ and (3.9), and let $a\in M^{\Phi}_{(\omega)}(\mathbf{R}^{2d})$. Then $\operatorname{Op}_{A,\varphi}(a)\in\mathscr{I}_{\Phi}(\omega_{1},\omega_{2})$, and
$\|\operatorname{Op}_{A,\varphi}(a)\|_{\mathscr{I}_{\Phi}(\omega_{1},\omega_{2})}$ $\leq Cd\hskip-0.80002pt\bar{}\hskip 1.99997pt^{-1}\exp(\|\varphi\|_{M^{\infty,1}_{(v)}})\|a\|_{M^{\Phi}_{(\omega)}},$ (3.17)
$a$ $\in M^{\Phi}_{(\omega)}(\mathbf{R}^{2d}),$
for some constant $C$ which is independent of $a$, $\varphi$ and $\Phi$.